### v2

# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

Abstract.
We define and study two kinds of directional expansiveness, weak and strong, for an action $T$ of $\mathbb{R}^{d}$ on a compact metric space $X$. We show that for $\mathbb{R}^{2}$ finite local complexity (FLC) tiling dynamical systems, weak and strong expansiveness are the same, and are both equivalent to a simple coding property. Then we show for the Penrose tiling dynamical system, which is FLC, there are exactly five non expansive directions, the directions perpendicular to the 5th roots of unity. We also study Raphael Robinson’s set of $24$ Penrose Wang tiles and show the corresponding Penrose Wang tile dynamical system is strictly ergodic. Finally, we study two deformations of the Penrose Wang tile system, one where the square Wang tiles are all deformed into a $52\pi/$ 5 rhombus, and another where they are deformed into a set of eleven tetragon tiles. We show both of these are topologically conjugate to the Penrose tiliing dynamical system.
Key words and phrases:
Penrose tiling, Wang tiles, tiling dynamical system, directional expansiveness.
2020 Mathematics Subject Classification:
37B52, 37B51, 37B05
1. Introduction
Let $X$ be a compact metric space with metric $\rho$. A homeomorphism $T$ of $X$ is called expansive (see [36]) if there exists a $\delta>0$ so that $\rho(T^{n}(x),T^{n}(y))<\delta$ for all $n\in\mathbb{N}$ implies $x=y$. Expansiveness means two distinct points cannot always stay close. Examples of expansive homeomorphisms include Anosov diffeomorphisms, Axiom A diffeomorphisms restricted to their non-wandering sets, and finite alphabet subshifts. It was shown in [27] and [22] that every expansive homeomorphism is a factor of a finite alphabet subshift, although the converse is false [8].
The definition of expansiveness readily generalizes to continuous $\mathbb{Z}^{d}$-actions $T$ on compact metric spaces $X$. However, when $d>1$, $\mathbb{Z}^{d}$-actions $T$ offer additional possibilities. Following in the footsteps of Milnor’s theory of directional entropy (see [23], [24]), Boyle and Lind [3] defined and studied directional expansiveness for $\mathbb{Z}^{d}$-actions. In this theory, an $e$-dimensional subspace ${\mathcal{V}}\subseteq\mathbb{R}^{d}$,$e<d$, is called an $e$-dimensional direction. Boyle and Lind [3] showed that for any $\mathbb{Z}^{d}$ action $T$, the set $\mathbb{E}_{e}(T)$ of $e$-dimensional expansive directions is open in the Grassmanian manifold ${\mathbb{G}}_{d,e}$, and the set $\mathbb{N}_{e}(T)$ of non-expansive directions is always non-empty. For every possible proper open set ${\mathcal{O}}\subseteq{\mathbb{G}}_{2,1}$ with #$\#({\mathcal{O}}^{c})>1$, except for the case #$\#({\mathcal{O}}^{c})=1$, Boyle and Lind [3] gave examples of $\mathbb{Z}^{2}$ actions $T$ with set $\mathbb{E}_{1}(T)={\mathcal{O}}$. Hochman [14] later provided examples with #$\#(\mathbb{N}_{1}(T))=1$.
A notion of expansiveness for flows $T$ (i.e., $\mathbb{R}$ actions) on compact metric spaces $X$ was introduced by Bowen and Walters in [4]. A flow $T$ is expansive if for all $\epsilon>0$, there is a $\delta>0$, so that for any $x,y\in X$ and any homeomorphism $h:\mathbb{R}\to\mathbb{R}$ with $h(0)=0$, $\rho(T^{t}x,T^{h(t)}y)<\delta$ for all $t\in\mathbb{R}$ implies $y=T^{s}x$ for some $|s|<\epsilon$. They proved that any fixed point free expansive flow on a 1-dimensional space $X$ is a quotient of a continuous suspension of a finite alphabet subshift. In the context of tiling dynamical systems $T$, Frank and Sadun [10] studied two types of expansiveness for $\mathbb{R}^{d}$-actions, which we call here weak and strong expansiveness. Strong expansiveness implies weak expansiveness, but not conversely in general. However, Frank and Sadun [10] showed that the converse holds for tiling dynamical systems satisfying the finite local complexity (FLC) property, (Definition 3.1).
In this paper, we define directional versions of both weak and strong expansiveness for an $\mathbb{R}^{d}$ action $T$ on a compact metric space $X$. The definition, which requires some care, is given in Section 2. And while weak directional expansiveness does not generally imply strong directional expansiveness, we show they are equivalent for finite local complexity (FLC) tiling dynamical systems. Also, in the FLC case, directional expansiveness is equivalent to a natural coding property.
Once the basic theory of directional expansiveness for $\mathbb{R}^{d}$-actions has been established, we move to our main study: describing the expansive directions for the Penrose tiling dynamical system. Our main theorem is the following:
Theorem 1.1.
For the Penrose tiling dynamical system $(\mathsf{X},\mathbb{R}^{2},T)$, (which is FLC), ((((see Section 4.1)))), there are exactly five $1$-dimensional non-expansive directions: the five directions perpendicular to $5^{th}$ roots of unity.
Our proof uses the translation equivariant version of de Bruijn’s [6] duality theory for pentagrid tilings, used in [29] to prove that the Penrose tiling dynamical system is an almost 1:1 extension of a Kronecker $\mathbb{R}^{2}$-action on the 4-torus $\mathbb{T}^{4}$ (stated here as Theorem 4.3). Our proof of Theorem 1.1 takes a closer look at factor map $\varphi$ in the proof of Theorem 4.3 in [29].
We also study Rapheel Robinson’s set of $24$ aperiodic Penrose Wang tiles (see [11]) and some related dynamical systems. In particular, we show the $\mathbb{Z}^{2}$ subshift of finite type $S$ corresponding to the Penrose Wang tiles is minimal, uniquely ergodic with pure point spectrum. We compute its point spectrum group and expansive directions.
1.1. Conventions
We will identify a $d$-tuple ${{\boldsymbol{w}}}=(w_{0},w_{1},\cdots,w_{d-1})\in\mathbb{R}^{d}$ with a $d\times 1$ column vector $[w_{0}w_{1}\cdots w_{d-1}]^{t}$. The usual dot product is denoted by ${\boldsymbol{w}}\cdot{\boldsymbol{v}}$. The vectors ${\boldsymbol{e}}_{0},{\boldsymbol{e}}_{1},\dots,{\boldsymbol{e}}_{d-1}$ denote the usual standard basis for $\mathbb{R}^{d}$. All norms are Euclidean, and for $d\geq 2$, $B_{r}({\boldsymbol{w}}):=\{{\boldsymbol{v}}\in\mathbb{R}^{d}:||{\boldsymbol{v}}-{\boldsymbol{w}}||<r\}$ denotes the open radius $r$ ball (or disc in the case $d=2$) in $\mathbb{R}^{d}$. For $d>1$, $\mathbb{T}^{d}:=\mathbb{R}^{d}/\mathbb{Z}^{d}$, with $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$ and $S^{1}=\{z\in\mathbb{C}:|z|=1\}$. We often identify $\mathbb{T}^{d}$ with $[0,1)^{d}$. For ${\boldsymbol{w}}\in\mathbb{R}^{d}$ we write $\lfloor{\boldsymbol{w}}\rfloor\in\mathbb{Z}^{d}$ for the coordinate-wise floor, and define $\{{\boldsymbol{w}}\}:={\boldsymbol{w}}-\lfloor{\boldsymbol{w}}\rfloor\in[0,1)^{d}$. Throughout, we let $\gamma:=\frac{1+\sqrt{5}}{2}\approx 1.6180\dots$ and $\alpha:=1/\gamma=\gamma-1\approx 0.6180\dots$.
1.2. Acknowledgements
Most of the results in this paper are from the first author’s 2021 Ph.D. dissertation [17] at the George Washington University. Theorem 1.1 answers a question Mike Boyle once asked the second author. The authors thank Sébastien Labbé for pointing out that Figure 10 was mislabeled in an earlier version.
2. Expansiveness and directional expansiveness
2.1. Group actions
Let $X$ be a compact metric space with metric $\rho$, and let $G$ be a locally compact complete metric abelian group. In this paper, the non-compact group $G$ will always be $\mathbb{Z}^{d}$, $\mathbb{R}^{d}$, or a subgroup of one of these, and the metric on $G$ will always be the Euclidean norm |||||$\cdot$||| | ⋅ | |.
A $G$-action $T$ on $X$ is a continuous mapping $T:G\times X\to X$, written $T^{\boldsymbol{g}}x:=T({\boldsymbol{g}},x)$, satisfying $T^{\boldsymbol{0}}={\rm id}$ and $T^{{\boldsymbol{g}}+{\boldsymbol{h}}}=T^{\boldsymbol{g}}\circ T^{\boldsymbol{h}}$ for all ${\boldsymbol{g}},{\boldsymbol{h}}\in G$. We assume the acting group $G$ is non-compact and think of a $G$-action as a dynamical system, denoted $(X,G,T)$ or sometimes $T$ for short. A $G$-action $T$ is free, if every $x\in X$ is aperiodic: $T^{\boldsymbol{g}}x=x$ implies ${\boldsymbol{g}}={{\boldsymbol{0}}}$, and locally free if there exists a $\omega>0$ so that $T^{\boldsymbol{g}}x=x$ implies $||{\boldsymbol{g}}||>\omega$.
A dynamical system $(X,G,T)$ is minimal if any closed $T$-invariant subset $K\subseteq X$ satisfies either $K$ or $K^{c}$ is empty, or equivalently, $X=\overline{O_{T}(x)}$ for some $x\in X$, where $O_{T}(x):=\{T^{\boldsymbol{g}}x:{\boldsymbol{g}}\in G\}$ is the orbit of $x$ under $T$, (see [7]). We say $(X,G,T)$ is uniquely ergodic if there is a unique $T$-invariant Borel probability measure $\mu$ on $X$. Strictly ergodic means both minimal and uniquely ergodic, either of which implies $T$ is free.
Definition 2.1.
For $(X,G,T)$ and $(Y,G,K)$, we say $K$ is a factor of $T$ (or $T$ is an extension of $K$) if there is a continuous surjection $\varphi:X\to Y$, the factor map, such that $K^{\boldsymbol{g}}(\varphi(x))=\varphi(T^{\boldsymbol{g}}x)$ for all ${\boldsymbol{g}}\in G$. An extension is almost $1\!\!:\!\!1$ if $Y_{1}:=\{x\in Y:\#(\varphi^{-1}(x))=1\}$ is dense $G_{\delta}$. In the case $\varphi$ is a homeomorphism, $K$ and $T$ are said to be topologically conjugate. We say $K$ and $T$ continuously orbit equivalent [^1] This is called “conjugacy of flows” in [4] but we avoid that terminology as it is too close to “topological conjugacy”. Sometimes this is called “flow equivalent”. if there is a homeomorphism $\psi:X\to Y$ so that $\psi(O_{T}(x))=O_{K}(\psi(x))$.
2.2. Kronecker dynamical systems
(See e.g., [30], [36]) The dual group $\widehat{G}$ of a locally compact metric abelian group $G$ (see e.g., [36]) is the locally compact metric abelian group of continuous homomorphisms ${\chi}:G\to S^{1}$ (in the compact-open topology). In particular, $\widehat{\mathbb{R}^{d}}=\mathbb{R}^{d}$ with $\chi_{\boldsymbol{h}}({\boldsymbol{t}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{t}})}$, and $\widehat{\mathbb{Z}^{d}}=\mathbb{T}^{d}$ with $\chi_{\boldsymbol{h}}({\boldsymbol{n}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{n}})}$, both uncountable. But the compact group $\mathbb{T}^{d}$ has a countable discrete dual $\widehat{\mathbb{T}^{d}}=\mathbb{Z}^{d}$ with $\chi_{\boldsymbol{n}}({\boldsymbol{h}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{n}})}$.
An eigenfunction $f$ for a dynamical system $T$, with eigenvalue $\chi\in\widehat{G}$, is a continuous222 It is also possible for some $T$ to have discontinuous eigenfunction, but for the examples we consider here, this is not the case. complex valued function satisfying $f(K^{\boldsymbol{g}}y)=\chi({\boldsymbol{g}})f(y)$. The set of eigenvalues $\Sigma_{T}\subseteq\widehat{G}$ is called the point spectrum of $T$. If $T$ is minimal, every eigenvalue $f$ is simple, $|f|=\text{const.}$., and $\Sigma_{T}\subseteq\widehat{G}$ is a countable discrete subgroup that is a topological conjugacy invariant.
For an infinite compact metric abelian group $Y$, let $i:G\to Y$ be a continuous injective homomorphism. We define an algebraic Kronecker dynamical system $(Y,G,K)$ to be the free $G$-action $K^{\boldsymbol{g}}y:=y+i({\boldsymbol{g}})$. It satisfies $\widehat{K}\cong\Sigma_{K}$ via $\chi\mapsto\chi\circ i$. Minimality is equivalent to $\overline{i(G)}=K$, or $\chi\mapsto\chi\circ i$ injective, (see [30]). More generally, we call $(X,G,T)$ Kronecker if it is topologically conjugate to an algebraic Kronecker dynamical system. For a free Kronecker system, the Halmos-von Neumann Theorem, (see e.g., [13]) says $\Sigma_{K}$ is a complete topological conjugacy invariant. Also, every countable dense subgroup $\widehat{Y}\subseteq\widehat{G}$ occurs as $\Sigma_{K}$ for some $K$.
Now consider $(X,G,T)$ and a Kronecker system $(Y,G,K)$ so that $\Sigma_{T}=\Sigma_{K}$. Then $(Y,G,K)$ is a factor of $(X,G,T)$, called its Kronecker (or maximal equicontinuous) factor. If the factor map $\varphi$ is almost 1:1 then we call $(X,G,T)$ almost automorphic, and call $\mathfrak{M}_{T}:=\{\#(\varphi^{-1}(y)):y\in Y\}\subseteq\mathbb{N}\cup\{+\infty\}$, the preimage multiplicities (called the thickness spectrum in [29], and called set of fiber cardinalities in [21]).
2.3. $\mathbb{Z}^{2}$ subshifts
For finite alphabet ${\mathcal{A}}=\{0,1,\dots,n-1\}$, $n\geq 2$, we call the product topology compact set ${\mathcal{A}}^{\mathbb{Z}^{2}}=\{z=(z_{\boldsymbol{m}})_{{\boldsymbol{m}}\in\mathbb{Z}^{2}}:z_{\boldsymbol{m}}\in{\mathcal{A}}\}$ the $\mathbb{Z}^{2}$ full shift space, equipped with the $\mathbb{Z}^{2}$ shift action $S$ defined by $(S^{\boldsymbol{n}}z)_{\boldsymbol{m}}=z_{{\boldsymbol{m}}-{\boldsymbol{n}}}$. The $\mathbb{Z}^{2}$ full shift dynamical system is $({\mathcal{A}}^{\mathbb{Z}^{2}},\mathbb{Z}^{2},S)$, and a $\mathbb{Z}^{2}$ subshift is $(Z,\mathbb{Z}^{2},S)$ where $Z\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$ is any closed and $S$-invariant.
Given a rectangle $B\subseteq\mathbb{Z}^{2}$, #$\#(B)>1$, a word on $B$ is $b=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}\in{\mathcal{A}}^{B}$. We say the word $b=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}$ occurs in $z\in{\mathcal{A}}^{\mathbb{Z}^{2}}$ if $(z_{{\boldsymbol{m}}-{\boldsymbol{n}}_{0}})_{{\boldsymbol{m}}\in B}=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}$ for some ${\boldsymbol{n}}_{0}\in\mathbb{Z}^{2}$. We call a subshift $Z$ a subshift of finite type (SFT) if there is a finite set ${\mathcal{F}}$ of forbidden words, so that now word in ${\mathcal{F}}$ occurs in any $z\in Z$. For a special case, called a $1$-step SFT, we let ${\mathcal{F}}={\mathcal{F}}_{0}\cup{\mathcal{F}}_{1}$ where ${\mathcal{F}}_{j}\subseteq{\mathcal{A}}^{\{{{\boldsymbol{0}}},{\boldsymbol{e}}_{j}\}}$. A $1$-step SFT can also be described in terms of permitted words ${\mathcal{P}}_{j}={\mathcal{A}}^{\{{{\boldsymbol{0}}},{\boldsymbol{e}}_{j}\}}\backslash{\mathcal{F}}_{j}$.
Let $(Z,\mathbb{Z},S)$ be a $\mathbb{Z}$ subshift, and let ${\mathcal{A}}\otimes{\mathcal{A}}:=\{a\otimes b:a,b\in{\mathcal{A}}\}$. We define the tensor product of $z,z^{\prime}\in Z$ by $z\otimes z\,^{\prime}:=(z_{n_{0}}\otimes z^{\prime}_{n_{1}})_{(n_{0},n_{1})\in\mathbb{Z}^{2}}\in({\mathcal{A}}\otimes{\mathcal{A}})^{\mathbb{Z}^{2}}$, and define the tensor square $Z\otimes Z\subseteq({\mathcal{A}}\otimes{\mathcal{A}})^{\mathbb{Z}^{2}}$ to be the set of all tensor product sequences $z\otimes z^{\prime}$ for $z,z^{\prime}\in Z$. The $\mathbb{Z}^{2}$ shift map on $Z\otimes Z$ is given by $(S\otimes S)^{(k_{0},k_{1})}(z_{n_{0}}\otimes z_{n_{1}})_{(n_{0},n_{1})\in\mathbb{Z}^{2}}=(z_{n_{0}-k_{0}}\otimes z_{n_{1}-k_{1}})_{(n_{0},n_{1})\in\mathbb{Z}^{2}}$. The tensor square dynamical system $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$ is a $\mathbb{Z}^{2}$ subshift, which is minimal, uniquely ergodic, or finite type, if and only if $(Z,\mathbb{Z},S)$ is the same.
2.4. Expansiveness for $\mathbb{Z}^{d}$ actions
A $\mathbb{Z}^{d}$ dynamical system $(X,\mathbb{Z}^{d},T)$ is expansive if there is a $\delta>0$ so that $\rho(T^{\boldsymbol{n}}x,T^{\boldsymbol{n}}y)<\delta$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{d}$ implies $x=y$. Expansiveness is an invariant of topological conjugacy (see [4] for the case $d=1$).
Now we want to describe expansiveness for an $\mathbb{R}^{d}$ dynamical system $(X,\mathbb{R}^{d},T)$, but simply replacing $\mathbb{Z}^{d}$ with $\mathbb{R}^{d}$ in the definition above does not quite work (see [4]). Actually, we will describe two versions of expansiveness for $\mathbb{R}^{d}$ dynamical systems, both of which were studied for $d=1$ by Bowen and Walters [4], (see also Frank and Sadun [10]). We call them weak and strong expansiveness.
Definition 2.2.
Let $(X,\mathbb{R}^{d},T)$ be a locally free $\mathbb{R}^{d}$ dynamical system. (i ) $T$ is called weakly expansive if for all $\epsilon>0$ there exists a $\delta>0$ so that for any $x,y\in X$, $\rho(T^{{\boldsymbol{t}}}x,T^{{\boldsymbol{t}}}y)<\delta$ for all ${{\boldsymbol{t}}}\in\mathbb{R}^{d}$ implies $y=T^{{{\boldsymbol{s}}}_{0}}x$ for some $||{{\boldsymbol{s}}}_{0}||<\epsilon$. (ii ) $T$ is called strongly expansive if for all $\epsilon>0$ there is a $\delta>0$ so that for any $x,y\in X$ and any homeomorphism $h:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $h({{\boldsymbol{0}}})={{\boldsymbol{0}}}$, $\rho(T^{{\boldsymbol{t}}}x,T^{{\boldsymbol{h}}({\boldsymbol{t}})}y)<\delta$ for all ${\boldsymbol{t}}\in\mathbb{R}^{d}$ implies $y=T^{{{\boldsymbol{s}}}_{0}}x$ for some $||{{\boldsymbol{s}}}_{0}||<\epsilon$.
Remarks 2.3.
  1. (1)
    Both weak and strong expansiveness are invariants of topological conjugacy.
  2. (2)
    The assumption “locally free” is a convenience that eliminates trivial cases like non isolated fixed points (see [4]). Locally free is automatic for FLC tiling dynamical systems (see Remark 3.7).
  3. (3)
    In the case $d=1$, Bowen and Walters [4] call our “strong expansiveness” (ii ) “expansiveness”. They also discuss our “weak expansiveness” (i ), without naming it, but dismiss it because it is not an invariant of “conjugacy of flows”: our “continuous orbit equivalence”.
  4. (4)
    Other versions of (i ) and (ii ) (a priori a little stronger) were studied for tiling dynamical systems by Frank and Sadun [10]. Their versions of (i ) and (ii ) , which are equivalent to ours and each other in the FLC case, have no $\epsilon$, and the conclusion is $||{\boldsymbol{s}}_{0}||<\delta$. They refer to (i ) as “geometric expansiveness”.
  5. (5)
    There are also versions of (i ) and (ii ) where the conclusion is simply that $y=T^{{\boldsymbol{s}}_{0}}x$, i.e., with no condition on $||{\boldsymbol{s}}_{0}||$. Gura [12] calls (i ) “separating” for $d=1$, and shows the horocycle flow satisfies (i ) but not (ii ) . For $d=1$, Katok and Hasselblatt [19] call this version of (ii ) for $d=1$ “expansiveness”. All (weak and strong) are equivalent in the FLC case.
2.5. Directional Expansiveness for $\mathbb{Z}^{d}$
For an $e$-dimensional direction (i.e., a subspace) ${\mathcal{V}}\subseteq\mathbb{R}^{d}$, $1\leq e<d$, we define ${\mathcal{V}}^{r}:=\bigcup_{{{\boldsymbol{v}}}\in{\mathcal{V}}}B_{r}({{\boldsymbol{v}}})$.
Definition 2.4 (Boyle and Lind, [3]).
Let $(X,\mathbb{Z}^{d},T)$ be a $\mathbb{Z}^{d}$ action and let ${\mathcal{V}}\subseteq\mathbb{R}^{d}$ be a $1\leq e<d$ dimensional direction. Then $T$ is expansive in the direction ${\mathcal{V}}$ if there exist a $\delta>0$ and an $r>0$ such that if $x,y\in X$ satisfy $\rho(T^{\boldsymbol{n}}x,T^{\boldsymbol{n}}y)\leq\delta$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{d}\cap{\mathcal{V}}^{r}$, then $x=y$.
In [3], the sets of $e$-dimensional expansive and non-expansive directions are denoted by $\mathbb{E}_{d}(T)$ and $\mathbb{N}_{d}(T)$. Let $(Z,\mathbb{Z}^{d},S)$ be a $\mathbb{Z}^{d}$ subshift, and let $J\subseteq\mathbb{R}^{d}$. We say a set $J$ determines $Z$ if for $z,z^{\prime}\in Z$, $z[J]=z^{\prime}[J]$ implies $z=z^{\prime}$. In the paragraph following (2–1) in [3], Boyle and Lind give the following useful characterization of directional expansiveness for a subshift.
Lemma 2.5 (Boyle and Lind, [3]).
For a $\mathbb{Z}^{d}$ subshift $(Z,\mathbb{Z}^{d},S)$ and an $e$-dimensional direction ${\mathcal{V}}\subseteq\mathbb{R}^{d}$, $1\leq e<d$, ${\mathcal{V}}$ is an expansive direction if and only if ${\mathcal{V}}^{r}$ determines $Z$ for some $r>0$.
Lemma 2.6.
Let $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$ be the tensor square of a $\mathbb{Z}$-subshift. For $J\subseteq\mathbb{Z}^{2}\subseteq\mathbb{R}^{2}$, suppose $\pi_{0}(J)=\mathbb{Z}$ and $\pi_{1}(J)=\mathbb{Z}$, where $\pi_{0},\pi_{1}:\mathbb{Z}^{2}\to\mathbb{Z}$ are the coordinate projections. Then $J$ determines $Z\otimes Z$.
Proof.
To find $z_{n_{0}}\otimes z^{\prime}_{n_{1}}$ for arbitrary $(n_{0},n_{1})\in\mathbb{Z}^{2}$ we use $\pi_{0}(J)=\mathbb{Z}$ and $\pi_{1}(J)=\mathbb{Z}$ to find $(n_{0},m_{1}),(m_{0},n_{1})\in J$. Then we use the first factor of $z_{n_{0}}\otimes z^{\prime}_{m_{1}}$, and the second factor of $z_{m_{0}}\otimes z^{\prime}_{n_{1}}$ to make $z_{n_{0}}\otimes z^{\prime}_{n_{1}}$. ∎
Corollary 2.7.
For any tensor square subshift $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$, the non-expansive directions $\mathbb{N}_{1}(S\otimes S)$ consist of exactly the horizontal and vertical directions.
Proof.
Let ${\mathcal{V}}$ be horizontal and let $r>0$. Find $z^{\prime},z^{\prime\prime}\in Z$ so that $z^{\prime}_{j}=z^{\prime\prime}_{j}$ for $|j|\leq r$ but $z^{\prime}_{k}\neq z^{\prime\prime}_{k}$ for some $|k|>r$. Then for any $z\in Z$, $(z\otimes z^{\prime})[{\mathcal{V}}^{r}]=(z\otimes z^{\prime\prime})[{\mathcal{V}}^{r}]$, since $z_{i}\otimes z^{\prime}_{j}=z_{i}\otimes z^{\prime\prime}_{j}$ for $i\in\mathbb{Z}$ and $|j|\leq r$, i.e., $(i,j)\in{\mathcal{V}}^{r}$. But $(z\otimes z^{\prime})\neq(z\otimes z^{\prime\prime})$, since $z^{\prime}_{k}\neq z^{\prime\prime}_{k}$ implies $z_{i}\otimes z^{\prime}_{k}\neq z_{i}\otimes z^{\prime\prime}_{k}$. Thus ${\mathcal{V}}^{r}$ does not determine $z\otimes z^{\prime}$ for any $r>0$. The same argument works for the vertical direction.
Now suppose ${\mathcal{V}}$ is neither vertical nor horizontal. Define $Q_{(n_{0},n_{1})}:=[n_{0}-\frac{1}{2},n_{0}+\frac{1}{2})\times[n_{1}-\frac{1}{2},n_{1}+\frac{1}{2})$. Choose $r>0$ large enough that $J:=\bigcup_{\{{\boldsymbol{n}}\in\mathbb{Z}^{2}:Q_{\boldsymbol{n}}\cap{\mathcal{V}}\neq\emptyset\}}Q_{\boldsymbol{n}}\subseteq{\mathcal{V}}^{r}$. $\pi_{0}(J)=\mathbb{Z}$ and $\pi_{1}(J)=\mathbb{Z}$, so the same holds for ${\mathcal{V}}^{r}$. Lemma 2.6 shows ${\mathcal{V}}^{r}$ determines $Z\otimes Z$, and the result follows from Lemma 2.5. ∎
2.6. Directional Expansiveness for $\mathbb{R}^{d}$
It turns out that the correct definition of direction ${\mathcal{V}}$ expansiveness for an $\mathbb{R}^{d}$ action $T$ is not simply the definition of expansiveness for $T$, restricted to ${\mathcal{V}}$, even though such a definition does work in some directional theories. The restriction of $T$ to ${\mathcal{V}}$, defined by $(T|_{\mathcal{V}})^{\boldsymbol{t}}:=T^{\boldsymbol{t}}$ for ${\boldsymbol{t}}\in{\mathcal{V}}$, is a ${\mathcal{V}}$ action, but can be made into an $\mathbb{R}^{e}$ action by a choice of a basis (see e.g., [3]). This does not affect questions of expansiveness.
Proposition 2.8.
Let $(X,\mathbb{R}^{d},T)$ be a locally free $\mathbb{R}^{d}$ action. Then for any direction ${\mathcal{V}}$, the restriction $(X,{\mathcal{V}},T|_{\mathcal{V}})$ to ${\mathcal{V}}$ is not weakly or strongly expansive.
Proof.
Suppose $T|_{\mathcal{V}}$ is weakly expansive and choose $\epsilon<\omega/2$ where $\omega$ is the constant for locally free. Let ${\boldsymbol{p}}\in\mathbb{R}^{d}\backslash{\mathcal{V}}$ with $||{\boldsymbol{p}}||=1$, and define a continuous real-valued function on $[0,1]\times X$ by $f(x,r):=\rho(x,T^{r{\boldsymbol{p}}}x)$. Then $f(0,x)=0$, and $f(r,x)>0$ for $r<\omega/2$. For $\delta>0$, find $20<r<\omega/$ 2 so that $f(r,x)<\delta$ for all $x\in X$. Fix $x\in X$ and let $y=T^{r{\boldsymbol{p}}}x$. Then for ${\boldsymbol{t}}\in{\mathcal{V}}$, $\rho(T^{{\boldsymbol{t}}}x,T^{{\boldsymbol{t}}}y)=\rho(T^{{\boldsymbol{t}}}x,T^{r{\boldsymbol{p}}}T^{\boldsymbol{t}}x)=f(T^{{\boldsymbol{t}}}x,r)<\delta$. Now, $T|_{\mathcal{V}}$ weakly expansive would mean $y=T^{r{\boldsymbol{p}}}x=T^{{\boldsymbol{s}}_{0}}x$ for some ${\boldsymbol{s}}_{0}\in{\mathcal{V}}$ with $||{\boldsymbol{s}}_{0}||<\epsilon<\omega/2$. But then $T^{{\boldsymbol{s}}_{0}-r{\boldsymbol{p}}}x=x$. Since $||{\boldsymbol{s}}_{0}-r{\boldsymbol{p}}||\leq||{\boldsymbol{s}}_{0}||+r||{\boldsymbol{p}}||<\omega$, locally free implies ${\boldsymbol{s}}_{0}=r{\boldsymbol{p}}$, contradicting $r{\boldsymbol{p}}\notin{\mathcal{V}}$. ∎
Definition 2.9.
Let $(X,\mathbb{R}^{d},T)$ be a locally free $\mathbb{R}^{d}$ action, and let ${\mathcal{V}}\subseteq\mathbb{R}^{d}$ be an $e$-dimensional direction, $1\leq e<d$. (i) $T$ is called weakly (directionally) expansive in the direction ${\mathcal{V}}$ if for all $\epsilon>0$ there is a $\delta>0$ such that for all $x,y\in X$, $\rho(T^{{\boldsymbol{t}}}x,T^{{\boldsymbol{t}}}y)<\delta$, for all ${{\boldsymbol{t}}}\in{\mathcal{V}}$ implies $y=T^{{\boldsymbol{s}}_{0}}x$ for some ${\boldsymbol{s}}_{0}\in\mathbb{R}^{d}$ with $||{{\boldsymbol{s}}}_{0}||<\epsilon$. (ii) $T$ is called strongly (directionally) expansive in the direction ${\mathcal{V}}$ if for all $\epsilon>0$ there is a $\delta>0$ such that for any $x,y\in X$ and any $h:{\mathcal{V}}\to{\mathcal{V}}$ with $h({{\boldsymbol{0}}})={{\boldsymbol{0}}}$, $\rho(T^{{\boldsymbol{t}}}x,T^{h({\boldsymbol{t}})}y)<\delta$ for all ${\boldsymbol{t}}\in{\mathcal{V}}$ implies $y=T^{{\boldsymbol{s}}_{0}}x$ for some ${\boldsymbol{s}}_{0}\in\mathbb{R}^{d}$ with $||{{\boldsymbol{s}}}_{0}||<\epsilon$.
3. Tiling dynamical systems
In this section we define finite local complexity (FLC) $\mathbb{R}^{2}$ tiling spaces and $\mathbb{R}^{2}$ tiling dynamical systems, and describe their basic properties (see e.g., [28] or [1]). Similar considerations hold for $\mathbb{R}^{d}$, $d>2$, but for simplicity we do not discuss them here.
3.1. Basic definitions
A tile is a subset $D\subseteq\mathbb{R}^{2}$ homeomorphic to a closed disk, with a Lebesgue measure zero boundary (e.g., a convex polygon). A tiling $x$ of $\mathbb{R}^{2}$ is a set of the tiles that cover $\mathbb{R}^{2}$, intersecting only on their boundaries. A tiling patch $x$ in $\mathbb{R}^{2}$ is a finite collection of the tiles whose union is a closed topological disc, and a tiling swatch $x$ is a collection of tiles whose union is closed and simply connected.
If $x$ is a tiling of $\mathbb{R}^{2}$, and $S\subseteq\mathbb{R}^{2}$, we let $x[S]$ be the smallest patch or swatch, the union of whose tiles contains $S$. Two tiles, patches, swatches or tilings are said to be equivalent if they are congruent via a translation. Sometimes geometrically equivalent tiles are “decorated” with colors or markings that make them inequivalent. We denote the translation of a tiling (or tile or patch or swatch) $x$ by ${{\boldsymbol{t}}}\in\mathbb{R}^{2}$ by $T^{{\boldsymbol{t}}}x:=\{D-{{\boldsymbol{t}}}:D\in x\}$. We call a set $p$ of inequivalent tiles in $\mathbb{R}^{2}$ a set of prototiles. A tiling (or path or swatch) $x$ by $p$ means $x$ each tile in $x$ is equivalent to a tile in $p$. We want to look at sets of all tilings by $p$, but without further restrictions, such a set is often too large.
Definition 3.1.
Let $p$ be a finite set of prototiles in $\mathbb{R}^{2}$. A local rule is a finite set $p^{(2)}$ of inequivalent $2$-tile patches by $p$. We let $X_{p}$ (or $p^{*}$) denote the set of all tilings (or patches) by $p$ such that every $2$-tile patch in any $x\in X_{p}$ (or any $y\in p^{*}$) is equivalent to a patch in $p^{(2)}$. We call $X_{p}$ the finite local complexity (abbreviated FLC) full tiling space defined by $p$, (with the local rule $p^{(2)}$ being implicit).
We will always assume that $p$ and $p^{(2)}$ are such that $X_{p}\not=\emptyset$. This is a nontrivial assumption in general (see Section 6.1), but is easy to ascertain in some cases (see Example 3.6), or is known in some others (see Remark 4.4).
There is a well known compact metric topology on any FLC tiling spaces $X_{p}$ in which two tilings are close if, after a small translation, they agree on a large ball around the origin. Here is one version of such a metric $\rho$ (see [28], [34]).
Definition 3.2.
For each tile $D\in p$, let $\delta(D)=\inf\{||{\boldsymbol{r}}||:T^{\boldsymbol{r}}D\cap D=\emptyset\}$ and let $\Delta_{p}:=\min_{D\in p}\delta(D)$. For $x,y\in X_{p}$, $x\neq y$, let $R(x,y):=\sup\left\{r:(T^{{\boldsymbol{r}}_{x}}x)[B_{r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{r}}_{y}}y)[B_{r}({{\boldsymbol{0}}})]\text{ some }||{\boldsymbol{r}}_{x}||,||{\boldsymbol{r}}_{y}||<1/r\right\}$, Define the tiling metric $\rho(x,y):=\min(\Delta_{p},1/R(x,y))$.
A well known result of Rudolph [33] is that an FLC full tiling space $X_{p}$ is a compact metric space (see also [28], [34]). The following facts will be useful.
Lemma 3.3.
Let $X_{p}$ be an FLC tiling space, and let $x,y\in X_{p}$.
  1. (1)
    If $(T^{\boldsymbol{r}}x)[\{{\boldsymbol{t}}\}]=(T^{{\boldsymbol{s}}}x)[\{{\boldsymbol{t}}\}]$ with $||{\boldsymbol{r}}-{\boldsymbol{s}}||<\Delta_{p}$ then ${\boldsymbol{r}}={\boldsymbol{s}}$.
  2. (2)
    If $\rho(x,y)<\delta<\min(1/3,\Delta_{p})$ and $r=1/\delta$, then there exists a ${\boldsymbol{t}}\in\mathbb{R}^{d}$ with $||{\boldsymbol{t}}||<\delta$ so that $x[B_{(8/9)r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{t}}}y)[B_{(8/9)r}({{\boldsymbol{0}}})]$.
  3. (3)
    If $\rho(T^{\boldsymbol{r}}x,T^{\boldsymbol{s}}x)<\Delta_{p}$ or $||{\boldsymbol{r}}-{\boldsymbol{s}}||<\Delta_{p}$ then $\rho(T^{\boldsymbol{r}}x,T^{\boldsymbol{s}}x)=||{\boldsymbol{r}}-{\boldsymbol{s}}||$.
Proof.
  1. (1)
    Suppose $x[\{{\boldsymbol{t}}\}]=(T^{{\boldsymbol{r}}}x)[\{{\boldsymbol{t}}\}]$ for $||{\boldsymbol{r}}||<\Delta_{p}$. Then for some $D\in x[\{{\boldsymbol{t}}\}]$ we also have $T^{\boldsymbol{r}}D\in x[\{{\boldsymbol{t}}\}]$, with $T^{\boldsymbol{r}}D\cap D\neq\emptyset$. So $T^{\boldsymbol{r}}D=D$ and ${\boldsymbol{r}}={{\boldsymbol{0}}}$.
  2. (2)
    If $0<\rho(x,y)=\delta=:1/r<\Delta_{p}$ then $(T^{{\boldsymbol{r}}}x)[B_{r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{s}}}y)[B_{r}({{\boldsymbol{0}}})]$ for some $||{\boldsymbol{r}}||,||{\boldsymbol{s}}||<\delta/2$. So $x[B_{r}({{\boldsymbol{r}}})]=(T^{{\boldsymbol{s}}-{\boldsymbol{r}}}y)[B_{r}({{\boldsymbol{r}}})]$ and ${\boldsymbol{t}}:={\boldsymbol{s}}-{\boldsymbol{r}}$ satisfies $||{\boldsymbol{t}}||<\delta$. It follows that $x[B_{(9/10)r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{t}}}y)[B_{(9/10)r}({{\boldsymbol{0}}})]$, because $31/r=\delta<1/$ 3 implies $B_{(9/10)r}({{\boldsymbol{0}}})\subseteq B_{r}({{\boldsymbol{r}}})$.
  3. (3)
    It suffices to prove $\rho(x,T^{\boldsymbol{r}}x)=||{\boldsymbol{r}}||$. Clearly $\rho(x,T^{\boldsymbol{r}}x)\leq||{\boldsymbol{r}}||$. Now suppose $\rho(x,T^{\boldsymbol{r}}x)=\delta=1/r<||{\boldsymbol{r}}||$. By (2) there is $||{\boldsymbol{s}}||<\delta$ with $(T^{\boldsymbol{s}}x)[B_{(8/9)r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{r}}}x)[B_{(8/9)r}({{\boldsymbol{0}}})]$, so by (1), ${\boldsymbol{r}}={\boldsymbol{s}}$, contradicting $\delta=||{\boldsymbol{s}}||<||{\boldsymbol{r}}||$. The converse is clear.
Definition 3.4.
For a full FLC tiling space $X_{p}$, we call any closed $T$-invariant subset $X\subseteq X_{p}$ an FLC tiling space (i.e., subspace) and call $(X,\mathbb{R}^{2},T)$ an FLC tiling dynamical system.
Let $X$ and $X^{\prime}$ be tiling spaces. For $r>0$, consider the sets of patches $X_{\{r\}}:=\{x[B_{r}({\boldsymbol{0}})]:x\in X\}$ and $X^{\prime}_{\{0\}}:=\{x^{\prime}[\{{\boldsymbol{0}}\}]:x^{\prime}\in X^{\prime}\}$.
Definition 3.5.
A map $C:X\to X^{\prime}$ is called local with local radius $r$ (see [28], [1]) if there is a map $c:X_{\{r\}}\to X^{\prime}_{\{0\}}$ so that $C(x)[\{{\boldsymbol{0}}\}]=c(x[B_{r}({\boldsymbol{0}})])$ for all $x\in X$.
Note that a local map is uniformly continuous, and satisfies $C(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}C(x)$ for $x\in X$ and ${\boldsymbol{t}}\in\mathbb{R}^{2}$. A local map is a factor map if it is onto. If a local homeomorphism $C:X\to X^{\prime}$ has a local inverse, we say $X$ and $X^{\prime}$ are mutually locally derivable (MLD), and say $C$ is a MLD topological conjugacy. Local maps between FLC tiling spaces are analogous to sliding block codes between $\mathbb{Z}^{d}$ subshifts. In the subshift case, the Curtis/Lyndon/Hedlund Theorem says every topological conjugacy is a sliding block code. However, for FLC tiling dynamical systems there are topological conjugacies that are not MLD (see [26], [15], [1], and Theorem 6.11).
Example 3.6.
For $j=0,1,\dots,4$, let ${{\boldsymbol{v}}}_{j}:=({\rm Re}(e^{2\pi ij/5}),{\rm Im}(e^{2\pi ij/5}))$, the $j$th $5$th root of unity, and let ${{\boldsymbol{v}}}^{\prime}_{j}:=(2/5){{\boldsymbol{v}}}_{j}$. We denote the prototile set by $p_{5}:=\{{{\boldsymbol{v}}}^{\prime}_{i}\wedge{{\boldsymbol{v}}}^{\prime}_{j}:0\leq i<j\leq 4\}$, where ${{\boldsymbol{v}}}^{\prime}_{i}\wedge{{\boldsymbol{v}}}^{\prime}_{j}:=\{t_{0}{{\boldsymbol{v}}}^{\prime}_{i}+t_{1}{{\boldsymbol{v}}}^{\prime}_{j}:t_{0},t_{1}\in[0,1]\}$. Then $p_{5}$ consists of $10$ rhombic tiles with edge length $2/5$, five with acute angle $52\pi/$ 5 and five with acute angle $102\pi/$ 10. Figure 1 (a) shows ${{\boldsymbol{v}}}^{\prime}_{0}\wedge{{\boldsymbol{v}}}^{\prime}_{4}$ and ${{\boldsymbol{v}}}^{\prime}_{0}\wedge{{\boldsymbol{v}}}^{\prime}_{3}$. All other tiles in $p_{5}$ are rotations of one of these. We take local rule $p^{(2)}_{5}$ to be the requirement that all tiles meet edge to edge. The corresponding FLC full tiling space satisfies $X_{p_{5}}\neq\emptyset$ since it contains periodic tilings (see Figure 1, (b)).
Remark 3.7.
The existence of periodic tilings $x\in X_{p_{5}}$ means that the FLC tiling dynamical system $(X_{p_{5}},\mathbb{R}^{2},T)$ is not minimal, uniquely ergodic, or even free. However, every FLC tiling dynamical system is locally free by (1) of Lemma 3.3.
Figure 1. (a) Two unmarked rhombic tiles, thick: ${\boldsymbol{v}}^{\prime}_{0}\wedge{\boldsymbol{v}}^{\prime}_{1}$ and thin ${\boldsymbol{v}}^{\prime}_{0}\wedge{\boldsymbol{v}}^{\prime}_{3}$. (b) A swatch of a periodic tiling $x\in X_{p_{5}}$, showing $X_{p_{5}}\neq\emptyset$, but also that $p_{5}$ is not an aperiodic prototile set.
3.2. Expansiveness and directional expansiveness for tiling dynamical systems
Theorem 3.8 (Frank-Sadun, [10]).
For $(X,\mathbb{R}^{2},T)$ an FLC tiling dynamical system, weak expansiveness is equivalent to strong expansiveness (we just say “expansive” in the FLC case).
Proposition 3.9.
Any FLC tiling dynamical system $(X,\mathbb{R}^{2},T)$ is expansive.
Proof.
Given $\epsilon>0$, let $\delta<\min\left(\epsilon,1/3,\Delta_{p}/2\right)$, $r=1/\delta$. Let $B_{0}=B_{8r/9}({{\boldsymbol{0}}})$. For a sequence ${\boldsymbol{t}}_{n}\in\mathbb{R}^{2}$, $n=1,2,\dots$, let $B_{n}:=B_{8r/9}(-{\boldsymbol{t}}_{n})$. Choose the ${\boldsymbol{t}}_{n}$ so that $\cup_{n\geq 1}B_{n}=\mathbb{R}^{2}$, and so that $I_{n}:=B_{n+1}\cap B_{n}\neq\emptyset$. Now suppose $x,y\in X_{p}$ satisfy $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)<\delta$ for all ${\boldsymbol{t}}\in\mathbb{R}^{2}$. Then for each $n=1,2,\dots$, $\rho(T^{{\boldsymbol{t}}_{n}}x,T^{{\boldsymbol{t}}_{n}}y)<\delta$, so that by (2) of Lemma 3.3, there exists $||{\boldsymbol{r}}_{n}||<\delta$ so that $(T^{{\boldsymbol{t}}_{n}}x)[B_{0}]=(T^{{\boldsymbol{t}}_{n}+{\boldsymbol{r}}_{n}}y)[B_{0}]$. Equivalently, $x[B_{n}]=(T^{{\boldsymbol{r}}_{n}}y)[B_{n}]$. We claim ${\boldsymbol{r}}_{n}={\boldsymbol{r}}_{1}$ for all $n\in\mathbb{N}$. Thus $x[B_{n}]=(T^{{\boldsymbol{r}}_{1}}y)[B_{n}]$ for all $n$, and since the $\mathbb{R}^{2}=\cup_{n\geq 1}B_{n}$, $y=T^{{\boldsymbol{r}}_{1}}x$. By (3) of Lemma 3.3, $||{\boldsymbol{r}}_{1}||=\rho(x,y)<\delta<\epsilon$. To prove the claim, we prove ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$ for all $n$. Let $C_{n}:=B_{n}\cap B_{n+1}$. Then $(T^{{\boldsymbol{r}}_{n}}y)[C_{n}]=x[C_{n}]=(T^{{\boldsymbol{r}}_{n+1}}y)[C_{n}]$. Since $||{\boldsymbol{r}}_{n+1}-{\boldsymbol{r}}_{n}||<2\delta<\Delta_{p}$, (1) of Lemma 3.3 implies ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$. ∎
We now discuss directional expansiveness for FLC tiling dynamical systems $(X,\mathbb{R}^{2},T)$. Any $1$-dimensional direction ${\mathcal{V}}$ in $\mathbb{R}^{2}$ is a line. We typically describe it either as the line generated by a vector ${\boldsymbol{v}}\neq{\boldsymbol{0}}$, or by its slope $m\in\mathbb{R}$, (with $m=\infty$ being a vertical line).
Proposition 3.10.
If $(X,\mathbb{R}^{2},T)$ is an FLC tiling dynamical system, then a direction ${\mathcal{V}}$ is weakly expansive if and only if it is strongly expansive in the direction ${\mathcal{V}}$.
Because of Proposition 3.10, we will from now on simply refer to expansive and non-expansive directions when discussing FLC tiling dynamical systems. Our proof of Proposition 3.10, closely follows the proof of Theorem 3.8 in [10]. We start with a lemma.
Lemma 3.11.
Let $(X,\mathbb{R}^{2},T)$, $X\subseteq X_{p}$, be an FLC tiling dynamical system, and let $0<2\delta<\min(1/3,\Delta_{p})$. For a $1$-dimensional direction ${\mathcal{V}}\subseteq\mathbb{R}^{2}$, and $x,y\in X$, let $h:{\mathcal{V}}\to{\mathcal{V}}$ be a homeomorphism such that $h({\boldsymbol{0}})={\boldsymbol{0}}$ and $\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)<\delta/3$ for all ${\boldsymbol{t}}\in{\mathcal{V}}$. Then $||h({\boldsymbol{t}})-{\boldsymbol{t}}||<\delta$ for all ${\boldsymbol{t}}\in\mathbb{R}$.
Proof.
We prove this by induction. Let $r=\delta/3$. Suppose there is a ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<r$ such that $||h({\boldsymbol{t}})-{\boldsymbol{t}}||\geq\delta$. By continuity we can find ${\boldsymbol{t}}_{0}\in{\mathcal{V}}$, $||{\boldsymbol{t}}_{0}||<r$, so that $||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=\delta$. Now, since $\rho(x,y)=\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)|_{{\boldsymbol{t}}={\boldsymbol{0}}}<\delta/3$, part (2) of Lemma 3.3 says there is ${\boldsymbol{r}}\in\mathbb{R}^{2}$ with $||{\boldsymbol{r}}||<\delta/3$ so that $x[B_{2r}({{\boldsymbol{0}}})]=(T^{\boldsymbol{r}}y)[B_{2r}({{\boldsymbol{0}}})]$. Since ${\boldsymbol{t}}_{0}\in B_{2r}({{\boldsymbol{0}}})$, we have $x[\{{\boldsymbol{t}}_{0}\}]=(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]$. Again, using part (2) of Lemma 3.3, $\rho(T^{{\boldsymbol{t}}_{0}}x,T^{h({\boldsymbol{t}}_{0})}y)<\delta/3$ implies $(T^{{\boldsymbol{t}}_{0}}x)[B_{2r}({{\boldsymbol{0}}})]=(T^{h({\boldsymbol{t}}_{0})+{\boldsymbol{s}}}y)[B_{2r}({{\boldsymbol{0}}})]$ for some $||{\boldsymbol{s}}||<\delta/3$. Thus $x[B_{2r}({{\boldsymbol{t}}_{0}})]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{s}}}y)[B_{2r}({\boldsymbol{t}}_{0})]$, which implies $x[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{s}}}y)[\{{\boldsymbol{t}}_{0}\}]$. It follows that $(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{s}}}y)[\{{\boldsymbol{t}}_{0}\}]$. Since $||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{s}}-{\boldsymbol{r}}||\leq||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||+||{\boldsymbol{r}}||+||{\boldsymbol{s}}||<\delta+\delta/3+\delta/3=2\delta<\Delta_{p}$, part (1) of Lemma 3.3 implies $h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}={\boldsymbol{s}}-{\boldsymbol{r}}$ which implies $\delta=||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=||{\boldsymbol{s}}-{\boldsymbol{r}}||<\delta/3+\delta/3<\delta$, a contradiction. This is the base of the induction.
For the induction step, we assume that, for some $k\geq 1$, no ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<rk$ has $||h({\boldsymbol{t}})-{\boldsymbol{t}}||\geq 2\delta/3$, and show the same for ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<(k+1)r$. Suppose, to the contrary, that some such ${\boldsymbol{t}}={\boldsymbol{t}}_{0}$ satisfies $||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=2\delta/3$. Let ${\boldsymbol{s}}_{0}=kt/(k+1)$ so that ${\boldsymbol{s}}_{0}\in{\mathcal{V}}$, $||{\boldsymbol{s}}_{0}||<kr$ and $||{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}||<r$. By the assumptions, $\rho(T^{{\boldsymbol{s}}_{0}}x,T^{h({\boldsymbol{s}}_{0})}y)<\delta/3$, and by induction, $||h({\boldsymbol{s}}_{0})-{\boldsymbol{s}}_{0}||<2\delta/3$, so part (2) of Lemma 3.3 implies
$\rho(T^{{\boldsymbol{s}}_{0}}x,T^{{\boldsymbol{s}}_{0}}y)\leq\rho(T^{{\boldsymbol{s}}_{0}}x,T^{h({\boldsymbol{s}}_{0})}y)+\rho(T^{h({\boldsymbol{s}}_{0})}y,T^{{\boldsymbol{s}}_{0}}y)<\delta/3+||h({\boldsymbol{s}}_{0})-{\boldsymbol{s}}_{0}||<\delta.$
Since $||{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}||<r$, the argument above shows $(T^{{\boldsymbol{s}}_{0}}x)[{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}]=(T^{{\boldsymbol{s}}_{0}+{\boldsymbol{r}}}y)[{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}]$ for some $||{\boldsymbol{r}}||<\delta/3$. Equivalently, $x[\{{\boldsymbol{t}}_{0}\}]=(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]$. But also, the assumption, $\rho(T^{{\boldsymbol{t}}_{0}}x,T^{h({\boldsymbol{t}}_{0})}y)<\delta/3$ implies $x[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{r}}^{\prime}}y)[\{{\boldsymbol{t}}_{0}\}]$ for some $||{\boldsymbol{r}}^{\prime}||<\delta/3$. So $(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{r}}^{\prime}}y)[\{{\boldsymbol{t}}_{0}\}]$. As above, part (1) of Lemma 3.3 implies $h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}={\boldsymbol{r}}^{\prime}-{\boldsymbol{r}}$, which yields the contradiction $32\delta/3=||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=||{\boldsymbol{r}}^{\prime}-{\boldsymbol{r}}||<2\delta/$ 3. ∎
Proof of Proposition 3.10.
Assume that $T$ is weakly expansive in the direction ${\mathcal{V}}$. Fix $\epsilon>0$ and find $\delta^{\prime}>0$ for weak expansiveness ((1) of Definition 2.9), satisfying $\delta^{\prime}<\Delta_{p}$. Let $x,y\in X$ and let $h=h_{x,y}$ be as in part (2) of Definition 2.9 satisfying: (i) $\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)<\delta^{\prime}/3$. Then by part (3) of Lemma 3.3 and Lemma 3.11 we have: (ii) $\rho(T^{\boldsymbol{t}}y,T^{h({\boldsymbol{t}})}y)=||h({\boldsymbol{t}})-{\boldsymbol{t}}||<2\delta^{\prime}/3$. Thus $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)\leq\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)+\rho(T^{h({\boldsymbol{t}})}y,T^{\boldsymbol{t}}y)<\delta^{\prime}/3+2\delta^{\prime}/3=\delta^{\prime}$, by (i) and (ii). It follows from weak expansiveness (part (1) of Definition 2.9) that $y=T^{{\boldsymbol{s}}_{0}}x$ for some ${\boldsymbol{s}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{s}}_{0}||<\epsilon$. This implies $T$ is strongly expansive (for $\delta=\delta^{\prime}/3$) in the direction ${\mathcal{V}}$. ∎
Definition 3.12.
Let ${\mathcal{V}}$ be a 1-dimensional direction. We say an FLC tiling dynamical system $T$ is direction ${\mathcal{V}}$ symbolically expansive if there is an $r>0$ so that $x[{\mathcal{V}}^{r}]=y[{\mathcal{V}}^{r}]$ for all $x,y\in X$ implies $x=y$.
The following should be compared to Lemma 2.5.
Proposition 3.13.
For an FLC tiling dynamical system $(X,\mathbb{R}^{2},T)$, direction ${\mathcal{V}}$ symbolic expansiveness is equivalent to direction ${\mathcal{V}}$ expansiveness.
Proof.
Suppose Definition 3.12 holds for $T$ in the direction ${\mathcal{V}}=\{t{\boldsymbol{v}}:t\in\mathbb{R}\}$, $||{\boldsymbol{v}}||=1$, for $r>0$. Fix $\epsilon>0$, and for $r^{\prime}=(2/3)\sqrt{3}r$, let $\delta<\min(\epsilon,1/r^{\prime},\Delta_{p}/2)$. Let ${\boldsymbol{t}}_{n}=nr^{\prime}{\boldsymbol{v}}$, $n\in\mathbb{Z}$. Then ${\mathcal{V}}^{r}\subseteq\cup_{n\in\mathbb{Z}}B_{r^{\prime}}({\boldsymbol{t}}_{n})$. Let $C_{n}:=B_{r^{\prime}}({\boldsymbol{t}}_{n})\cap B_{r^{\prime}}({\boldsymbol{t}}_{n+1})$. Now suppose $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)<\delta$ for all ${\boldsymbol{t}}\in{\mathcal{V}}$. For each $n\in\mathbb{Z}$ there is a small shift, $||{\boldsymbol{r}}_{n}||<\delta$, such that $(T^{{\boldsymbol{t}}_{n}}x)[B_{r^{\prime}}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{r}}_{n}}T^{{\boldsymbol{t}}_{n}}y)[B_{r^{\prime}}({{\boldsymbol{0}}})]$, or equivalently $x[B_{r^{\prime}}({-{\boldsymbol{t}}_{n}})]=(T^{{\boldsymbol{r}}_{n}}y)[B_{r^{\prime}}({-{\boldsymbol{t}}_{n}})]$. We claim ${\boldsymbol{r}}_{n}={\boldsymbol{r}}_{n+1}$ for all $n\in\mathbb{Z}$. we have $(T^{{\boldsymbol{r}}_{n}}y)[C_{n}]=x[C_{n}]=(T^{{\boldsymbol{r}}_{n+1}}y)[C_{n}]$ where $||{\boldsymbol{r}}_{n}||,||{\boldsymbol{r}}_{n+1}||<\delta$. Since $||{\boldsymbol{r}}_{n+1}-{\boldsymbol{r}}_{n}||<\Delta_{p}$, (1) of Lemma 3.3 implies ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$. So $x[B_{r^{\prime}}({-{\boldsymbol{t}}_{n}})]=(T^{{\boldsymbol{r}}_{0}}y)[B_{r^{\prime}}({-{\boldsymbol{t}}_{n}})]$ for all $n$. But the definition of $r^{\prime}$ implies ${\mathcal{V}}^{r}\subseteq\cup_{n\in\mathbb{Z}}B_{r^{\prime}}({\boldsymbol{t}}_{n})$. Thus $x[{\mathcal{V}}^{r}]=(T^{{\boldsymbol{r}}_{0}}y)[{\mathcal{V}}^{r}]$, and by Definition 3.12, $x=T^{{\boldsymbol{r}}_{0}}y$ with $||{\boldsymbol{r}}_{0}||<\delta<\epsilon$.
Conversely, suppose $T$ is weakly expansive in direction ${\mathcal{V}}$. Given $\epsilon>0$, let $\delta<\min(\epsilon,\Delta_{p})$, and let $r>1/\delta$. Now suppose $x[{\mathcal{V}}^{r}]=y[{\mathcal{V}}^{r}]$. Then $x[B_{r}(-{\boldsymbol{t}})]=y[B_{r}(-{\boldsymbol{t}})]$ for all ${\boldsymbol{t}}\in{\mathcal{V}}$, or equivalently, $(T^{\boldsymbol{t}}x)[B_{r}({{\boldsymbol{0}}})]=(T^{\boldsymbol{t}}y)[B_{r}({{\boldsymbol{0}}})]$ for all ${\boldsymbol{t}}\in{\mathcal{V}}$. Thus $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)=1/r<\delta$. By weak expansiveness in direction ${\mathcal{V}}$, $y=T^{{\boldsymbol{s}}_{0}}x$ with $||{\boldsymbol{s}}_{0}||<\epsilon<\delta<\Delta_{p}$. But $(T^{{\boldsymbol{s}}_{0}}x)[{\mathcal{V}}^{r}]=x[{\mathcal{V}}^{r}]$, so by (1) of Lemma 3.3, we have ${\boldsymbol{s}}_{0}={{\boldsymbol{0}}}$ and $x=y$. ∎
4. The Penrose tiling dynamical system
4.1. Penrose tilings
In the next two sections (4 and 5), we discuss two types of Penrose tilings: the arrowed rhombic Penrose tilings and the unmarked rhombic Penrose tilings (see [28], [29], [1]). Later, in Section 6, we will discuss two additional types of tilings whose tiling dynamical systems are topologically conjugate to the Penrose tiling dynamical system, (see Theorem 6.11).
The arrowed rhombic Penrose tiles, denoted by $p_{5{\rm a}}$, are versions of the tiles from $p_{5}$ (Example 3.6) marked with arrows. Two of these tiles are shown in Figure 2(a). The entire set $p_{5{\rm a}}$ of $20$ prototiles is obtained by taking all rotations of these two tiles by multiples of $102\pi/$ 10. To make an FLC prototile set, the local rule $p^{(2)}_{5{\rm a}}$ requires two tiles to meet edge to edge with matching arrows (see Figure 2).
Figure 2. (a) two arrowed rhombic Penrose tiles, (b) arrowed rhombic Penrose tiling patch, and (c) the same patch with arrows erased. The erasing map $E$ has a local inverse $E^{-1}$.
The erasing map $E:X_{p_{\text{5{\rm a}}}}\to X_{p_{5}}$, which erases the arrows, is not onto (see [6]) since no periodic tiling can be correctly decorated with arrows. Clearly $E(X_{p_{\text{5{\rm a}}}})$ is a tiling space.
Definition 4.1.
The unmarked Penrose tiling tiling space is defined by $\mathsf{X}:=E(X_{p_{\text{5{\rm a}}}})\subseteq X_{p_{5}}$, and we call the $(\mathsf{X},\mathbb{R}^{2},T)$ the (unmarked rhombic) Penrose tiling dynamical system.
Lemma 4.2 (deBruijn, [6]).
The erasing map $E:X_{p_{\text{5{\rm a}}}}\to X_{p_{5}}$ is 1:1, and for $\mathsf{X}=E(X_{p_{\text{5{\rm a}}}})$, the inverse $E^{-1}:\mathsf{X}\to X_{p_{\text{5{\rm a}}}}$ is an MLD topological conjugacy.
Proof.
It is shown in [6] that, up to rotation there are only seven vertex configurations (see Figure 3) among all the $x\in X_{p_{\text{5{\rm a}}}}$. Types “S” and “S5” are the only two that are the same in $E(x)$, and since they can never be adjacent, they can be distinguished in $E(x)$ by the patterns on at least one adjacent vertex. ∎
Figure 3. Up to rotation there are seven arrowed vertex configurations in all $x\in X_{p_{\text{5{\rm a}}}}$. The labels for the types are from [6].
4.2. The Kronecker model for Penrose tilings
Let $\mathbb{T}^{5}_{0}=\{{\boldsymbol{u}}\in\mathbb{T}^{5}:\{u_{0}+\cdots+u_{4}\}=0\}$, a closed subtorus of $\mathbb{T}^{5}$, homeomorphic to $\mathbb{T}^{4}$. For ${\boldsymbol{v}}_{j}=({\rm Re}(e^{2\pi i/5}),{\rm Im}(e^{2\pi i/5}))$, the $j$th 5th root of unity, let $W=[{\boldsymbol{v}}_{0}^{t};{\boldsymbol{v}}_{1}^{t};\dots;{\boldsymbol{v}}_{4}^{t}]$ be the $5\times 2$ matrix rows ${\boldsymbol{v}}_{j}$. Let $V$ be the $4\times 2$ matrix obtained by deleting the row of ${\boldsymbol{v}}_{3}^{t}$ from $W$. For ${\boldsymbol{t}}\in\mathbb{R}^{2}$, define $K^{\boldsymbol{t}}{\boldsymbol{u}}=\{{\boldsymbol{u}}+W{\boldsymbol{t}}\}$ on $\mathbb{T}^{5}$ and $L^{{\boldsymbol{t}}}{{\boldsymbol{u}}}=\{{{\boldsymbol{u}}}+V{{\boldsymbol{t}}}\}$ on $\mathbb{T}^{4}$. Clearly $(\mathbb{T}^{5},\mathbb{R}^{2},K)$ and $(\mathbb{T}^{4},\mathbb{R}^{2},L)$ are Kronecker systems, with $\Sigma_{K}=\Sigma_{L}=V^{t}\mathbb{Z}^{5}=W^{t}\mathbb{Z}^{4}=\mathbb{Z}[e^{2\pi i/5}]\subseteq\mathbb{R}^{2}$, the decomplexification of the complex ring generated by $e^{2\pi i/5}$. Note that $K$ is not minimal since $\mathbb{T}^{5}_{0}$ is $K$ invariant, but $L$ is minimal since $W^{t}$ is injective on $\mathbb{Z}^{4}$. Moreover, $(\mathbb{T}^{4},\mathbb{R}^{2},L)$ is topologically conjugate to $(\mathbb{T}^{5}_{0},\mathbb{R}^{2},K|_{\mathbb{T}^{5}_{0}})$. This is the unique minimal Kronecker $\mathbb{R}^{2}$ dynamical system with $\Sigma_{K}=\mathbb{Z}[e^{2\pi i/5}]$. The following is the main result of [29].
Theorem 4.3 (see [29]).
The Penrose tiling dynamical system $(\mathsf{X},\mathbb{R}^{2},T)$ is strictly ergodic and almost automorphic, with $\Sigma_{T}=\mathbb{Z}[e^{2\pi i/5}]$ and $\mathfrak{M}_{T}=\{1,2,10\}$.
Previously, Penrose [25] had obtained the basic properties of Penrose tilings, (many of which also follow from Theorem 4.3), although not expressing his results in dynamical terminology. He showed that every $x\in X_{p_{\text{5{\rm a}}}}$ is aperiodic, so $T$ is free, and also that $T$ is minimal and uniquely ergodic (see [29]). Penrose’s proofs of these property are based on the following remarkable scaling property for Penrose tilings.
Proposition 4.4 (Penrose, [25], see also [11], [28]).
There is a MLD topological conjugacy (see Definition 3.5 and the comments that follow it) $C:X_{p_{\text{5}}}\to X_{(1/\gamma)p_{\text{5}}}$ (or $C:\mathsf{X}\to(1/\gamma)\mathsf{X}$ ), where $(1/\gamma)p_{\text{5{\rm a}}}$ (or $(1/\gamma)\mathsf{X}$) is same set of tilings, scaled down by $1/\gamma$, (see [29] for a picture of $C$).
In addition, the map $C^{\prime}:=\gamma C:X_{p_{\text{5{\rm a}}}}\to X_{p_{\text{5{\rm a}}}}$, or $C^{\prime}:\mathsf{X}\to\mathsf{X}$, (which scales back up to the original scale) called a tiling substitution or inflation, is also a topological conjugacy (see [28]). Note that in [1] this would be called an improper inflation as opposed to a stone inflation, because the map $C$ does not perfectly tile each $D\in p_{\text{5}}$ with tiles $D^{\prime}\in(1/\gamma)p_{\text{5{\rm a}}}$. The scaled map $c^{\prime}=\gamma c$ corresponding to $C^{\prime}:=\gamma C$ (see Remark 3.5) can be applied iteratively to tiling patches rather than tiles, allowing one to construct ever larger patches thereby proving $X_{p_{\text{5{\rm a}}}}\neq\emptyset$.
4.3. Grid tiling dynamical system
Here we recall some ideas from the proof of Theorem 4.3 from [29] that we will use in the proof of Theorem 1.1. That proof is based on deBruijn’s theory of grid tilings [6], (see also [28] or [1]).
[…]
$\ell_{j,k}=\ell_{j,k}(u_{j}):=\{{\boldsymbol{v}}_{j}\}^{\perp}+(-u_{j}+k){\boldsymbol{v}}_{j}\subseteq\mathbb{R}^{2}.$ (4.1)
We view the set $y({\boldsymbol{u}}):=\{\ell_{j,k}(u_{j}):j=0,1,\dots,4;k\in\mathbb{Z}\}$ of all grid lines as a non FLC tiling of $\mathbb{R}^{2}$ by convex polygonal grid tiles, (see Figure 4), which we call a grid tiling. We denote the set of all grid tilings by $Y:=\{y({\boldsymbol{u}}):{\boldsymbol{u}}\in\mathbb{T}^{5}_{0}\}$.
Figure 4. (a) Patch of nonsingular grid tiling $y({\boldsymbol{u}})\in Y_{1}$. Note that grid tiles are often almost vanishingly small. One grid patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ is marked in grey, and (b) also shown right. As described in Definition 5.3, this grid patch has symbol $\sigma(g_{\boldsymbol{n}}({\boldsymbol{u}}))=0\otimes 1$. The dual $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$ (see Lemma 5.8) of this grid patch, where $x=y({\boldsymbol{u}})^{*}$, is of type $w^{20}$. It is shown in Figure 9.
The mapping ${\boldsymbol{u}}\mapsto y({\boldsymbol{u}}):\mathbb{T}^{5}_{0}\to Y$ is a bijection, continuous in the topology333 The topology on $Y$ can also be defined geometrically using a non-FLC version of the tiling metric, (see [28] or [10]). that $\mathbb{T}^{5}_{0}$ pushes forward to $Y$. Thus $(Y,\mathbb{R}^{2},T)$ is a non FLC tiling dynamical system, where $T$ is the $\mathbb{R}^{2}$ translation action on $Y$.
Lemma 4.5.
The map $y:\mathbb{T}^{5}_{0}\to Y$ satisfies $y(K^{\boldsymbol{t}}{\boldsymbol{u}})=T^{\boldsymbol{t}}(y({\boldsymbol{u}}))$. Thus $(Y,\mathbb{R}^{2},T)$ is the unique minimal Kronecker dynamical system with spectrum $\Sigma_{K}=\mathbb{Z}[e^{2\pi i/5}]$.
Proof.
The $\mathbb{Z}^{d}$ valued function
${\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}}):=\lfloor W{\boldsymbol{s}}+{\boldsymbol{u}}\rfloor-\lfloor{\boldsymbol{u}}\rfloor,$ (4.2)
where ${\boldsymbol{s}}\in\mathbb{R}^{2}$, ${\boldsymbol{u}}\in\mathbb{R}^{5}$, satisfies ${\boldsymbol{m}}({{\boldsymbol{0}}},{\boldsymbol{u}})={{\boldsymbol{0}}}$, ${\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}}+{\boldsymbol{k}})={\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})$, ${\boldsymbol{k}}\in\mathbb{Z}^{5}$, so is well defined on $\mathbb{T}^{5}_{0}$. It also satisfies the cocycle equation ${\boldsymbol{m}}({\boldsymbol{t}}+{\boldsymbol{s}},{\boldsymbol{u}})-{\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})={\boldsymbol{m}}({\boldsymbol{t}},K^{\boldsymbol{s}}{\boldsymbol{u}})$. Fixing ${\boldsymbol{u}}\in\mathbb{T}^{5}_{0}$, we see that ${\boldsymbol{m}}({\boldsymbol{t}},{\boldsymbol{u}})$ is constant (interiors of) tiles $E\in y({\boldsymbol{u}})$, and takes a different value on each tile. The left side of the cocycle equation is piecewise constant on the tiles of $y({\boldsymbol{u}})-{\boldsymbol{s}}=T^{\boldsymbol{s}}y({\boldsymbol{u}})$, whereas the right side is piecewise constant on the tiles of $y(K^{\boldsymbol{s}}{\boldsymbol{u}})$. ∎
4.4. Duality and nonsingular tilings
For an edge to edge tiling $y$ of $\mathbb{R}^{2}$ by polygonal tiles, let $v(y)$ be its vertices, and $e(y)$ be its edges, with the tiles in $y$ being its faces. Another edge to edge tiling $y^{*}$ by polygonal tiles is said to be geometrically dual to $y$ if there are bijections $y\overset{*}{\leftrightarrow}v(x)$, $v(y)\overset{*}{\leftrightarrow}x$ and $e(y)\overset{*}{\leftrightarrow}e(x)$, such that $\ell^{*}\in e(x)$, $\ell\in e(y)$ implies $\ell^{*}\perp\ell$.
We will now describe a geometric duality (see [6], [29] for more details) between the grid tilings $y=y({\boldsymbol{u}})\in Y$ and a certain set of tilings $Y^{*}\supseteq X_{p5}$. First, for a tile $E\in y({\boldsymbol{u}})$, we define the vertex
$E^{*}:=(2/5)W^{t}({\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})+\{{\boldsymbol{u}}\})\in v(y^{*}),$ (4.3)
for arbitrary ${\boldsymbol{s}}\in{\rm int}(E)$. We connect two vertices $E_{1}^{*},E_{2}^{*}\in v(y^{*})$ by an edge $\ell^{*}\in e(y^{*})$ if their distance is $2/5$, or equivalently, if $E_{1},E_{2}\in y$ are adjacent across an edge $\ell\in e(y)$. The edges $e(y^{*})$ divide $\mathbb{R}^{2}$ into the polygonal tiles $D$ that make up $y^{*}$. In particular, if ${\boldsymbol{b}}\in v(y)$ is an $n$-fold crossing, then it is the common vertex of $2n$ tiles $E_{0},E_{1},\dots,E_{2n}\in y$, going clockwise. The duals $E^{*}_{0},E^{*}_{1},\dots,E^{*}_{2n}\in v(y^{*})$ and the edges in $e(y^{*})$ connecting them enclose a $2n$-gon ${\boldsymbol{b}}^{*}:=D\in y^{*}$. The case $n=2$ is shown in Figure 5.
Lemma 4.6 (See [29]).
For $y=y({\boldsymbol{u}})\in Y$, the map $y({\boldsymbol{u}})\mapsto y^{*}({\boldsymbol{u}})$ is injective and satisfies
$(T^{\boldsymbol{t}}y({\boldsymbol{u}}))^{*}=y(K^{\boldsymbol{t}}{\boldsymbol{u}})^{*}=T^{\boldsymbol{t}}(y({\boldsymbol{u}})^{*}).$ (4.4)
Proof.
We first prove (4.4). The first equality follows from Lemma 4.5. For the second, we have that $v(y({\boldsymbol{u}})^{*})=\frac{2}{5}W^{t}({\boldsymbol{m}}(\mathbb{R}^{2},{\boldsymbol{u}})-\{{\boldsymbol{u}}\})=\frac{2}{5}W^{t}(\lfloor W\mathbb{R}^{2}+{\boldsymbol{u}}\rfloor-{\boldsymbol{u}})$. Then $K^{\boldsymbol{t}}{\boldsymbol{u}}=W{\boldsymbol{t}}+{\boldsymbol{u}}$ and $\frac{2}{5}W^{t}W=I$ imply $v(y(K^{\boldsymbol{t}}{\boldsymbol{u}})^{*})=\frac{2}{5}W^{t}(\lfloor W\mathbb{R}^{2}+W{\boldsymbol{t}}+{\boldsymbol{u}}\rfloor-(W{\boldsymbol{t}}+{\boldsymbol{u}}))=\frac{2}{5}W^{t}(\lfloor W\mathbb{R}^{2}+{\boldsymbol{u}}\rfloor-{\boldsymbol{u}})-{\boldsymbol{t}}=v(y({\boldsymbol{u}})^{*})-{\boldsymbol{t}}$. Since $v(y^{*})$ determines $y^{*}$, we have $y^{*}(K^{\boldsymbol{t}}{\boldsymbol{u}})=y^{*}({\boldsymbol{u}})-{\boldsymbol{t}}=T^{\boldsymbol{t}}y^{*}({\boldsymbol{u}})$.
The duality is clearly injective up to equivalence: if $y_{1},y_{2}\in Y$ are not translates then $y^{*}_{1},y^{*}_{2}$ cannot be translates, [6]. The injectivity of $y\mapsto y^{*}$ then follows from (4.4). ∎
We call a grid tiling $y=y({\boldsymbol{u}})\in Y$ nonsingular if every vertex ${\boldsymbol{b}}\in v(y)$ is a 2-fold crossing. We denote the set of all nonsingular grid tilings by $Y_{1}$. Note that $y^{-1}(Y_{1})\subseteq\mathbb{T}^{5}_{0}$ is dense $G_{\delta}$, as is $Y_{1}\subseteq Y$. Now suppose $y=y({\boldsymbol{u}})\in Y_{1}$, and the vertex ${\boldsymbol{b}}\in v(y)$ is a ($2$-fold) crossing of $\ell_{i,k}$ and $\ell_{j,k^{\prime}}$, $i<j$. Then its dual ${\boldsymbol{b}}^{*}$ in $y^{*}$ is a type ${\boldsymbol{v}}^{\prime}_{i}\wedge{\boldsymbol{v}}^{\prime}_{j}$ tile in $p_{5}$ (see Figure 5). Thus $y\in Y_{1}$ implies $y^{*}\in X_{p_{5}}$. But in fact, much more is true.
Theorem 4.7 (deBruijn, [6]).
If $y=y({\boldsymbol{u}})\in Y_{1}$ then $x({\boldsymbol{u}}):=y({\boldsymbol{u}})^{*}\in\mathsf{X}$. In other words, the dual $x({\boldsymbol{u}})$ of any nonsingular grid tiling $y({\boldsymbol{u}})$ is an unmarked Penrose tiling.
Figure 5. A vertex ${\boldsymbol{b}}\approx{\boldsymbol{0}}\in v(y)$ (left) that is a 2-fold crossing of a 0-grid line and a 1-grid line. The values of ${\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})$ on the four tiles $E_{0},E_{1},E_{2}$ and $E_{3}$ adjacent to ${\boldsymbol{b}}$ have values ${\boldsymbol{m}}={{\boldsymbol{0}}}$, ${\boldsymbol{m}}=-{\boldsymbol{e}}_{1}$, ${\boldsymbol{m}}=-({\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1})$ and ${\boldsymbol{m}}=-{\boldsymbol{e}}_{1}$. Assuming $\{{\boldsymbol{u}}\}={\boldsymbol{0}}$, their duals (right) are $E_{0}^{*}={{\boldsymbol{0}}}$, $E_{1}^{*}=-{\boldsymbol{v}}^{\prime}_{0}$, $E_{2}^{*}=-({\boldsymbol{v}}^{\prime}_{0}+{\boldsymbol{v}}^{\prime}_{1})$, $E_{3}^{*}=-{\boldsymbol{v}}^{\prime}_{1}$, which are the vertices of a tile ${\boldsymbol{b}}^{*}$ (right), of type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}_{1}^{\prime}$.
Let $\mathsf{X}_{1}:=Y^{*}_{1}$, called the nonsingular Penrose tilings, and let $\varphi:\mathsf{X}_{1}\to Y_{1}$ be the inverse of the duality map.
Lemma 4.8 (See [29]).
The map $\varphi:\mathsf{X}_{1}\to Y_{1}$ is uniformly continuous and has a continuous extension to $\varphi:\mathsf{X}\to Y$ where $\mathsf{X}:=\overline{\mathsf{X}_{1}}$.
This is essentially deBruijn’s observation [6] that all Penrose tilings are limits of nonsingular Penrose tilings. The extension $\varphi:\mathsf{X}\to Y$ is the factor map in Theorem 4.3. The fact that $\varphi$ has a single preimage on the $G_{\delta}$ set $Y_{1}$ shows that $\varphi$ is almost 1:1.
4.5. The singular cases
Figure 6. (a) The $3$-fold crossings along the spine $\ell_{0}$ for $y({\boldsymbol{u}})\in Y_{2}$ and the unfilled work in its dual $y({\boldsymbol{u}})^{*}$. (b) and (c) show the $+$ and -- resolutions of the worm.
If $y({\boldsymbol{u}})\notin Y_{1}$ it means that some of crossings in $y({\boldsymbol{u}})$ are not $2$-fold. We call $y({\boldsymbol{u}})$ a singular grid tiling. Let $n$ be the highest crossing order in $y({\boldsymbol{u}})$. It is easy to see that ${\boldsymbol{u}}\in\mathbb{T}^{5}_{0}$ implies that $n=4$ never occurs (see [6]). Thus there are two cases: Case A, $n=3$, which we will see corresponds to $Y_{2}$ in Theorem 4.3, and Case B, $n=5$, which corresponds to $Y_{10}$.
Case A: Let ${\boldsymbol{b}}$ be one of the 3-fold crossings in $y({\boldsymbol{u}})$. As shown in [6] (see also [29]), there is a grid line $\ell_{j,k}$ through ${\boldsymbol{b}}$, we call it a spine, such that all the crossings along $\ell_{j,k}$ are $3$-fold, and all the other crossings in $y({\boldsymbol{u}})$ are $2$-fold. The $3$-fold crossings along $\ell_{j,k}$ are of two kinds: wide is a crossing of $\ell_{j}$ with (some) $\ell_{j+2,k^{\prime}}$ and $\ell_{j-2,k^{\prime\prime}}$ (where the arithmetic for $j$ is $\!\!\!\mod\!5$), and narrow is a crossing of $\ell_{j}$ with $\ell_{j+1,k^{\prime}}$ and $\ell_{j-1,k^{\prime\prime}}$, (see Figure 6(a)).
The duals of $3$-fold crossings are hexagons, respectively wide or narrow. The hexagons line up in $y({\boldsymbol{u}})^{*}$ in configuration called an unfilled worm, Figure 6(b), and we call the line $l$ along the center of the worm its axis. Since hexagons are not in $p_{5}$, the dual $y({\boldsymbol{u}})^{*}$ is not in $X_{p_{5}}$. We can obtain a tiling in $X_{p_{5}}$ by “filling” each hexagon in $y({\boldsymbol{u}})^{*}$ by tiling it with three tiles from $p_{5}$. This can be done in one of two directions: direction $+$, meaning a single tile sits on the side facing ${\boldsymbol{v}}_{j}$, and direction --, meaning two tiles sit on the side facing ${\boldsymbol{v}}_{j}$, (see Figure 6(b) and (c)).
To obtain a Penrose tiling, all the hexagons must be filled in the same direction [6]. We denote the two resulting Penrose tilings as $x^{+}({\boldsymbol{u}}),x^{-}({\boldsymbol{u}})\in\mathsf{X}$, Figure 6(b) and (c). The column of filled hexagons in $x^{+}({\boldsymbol{u}})$ or $x^{-}({\boldsymbol{u}})$ is called a filled worm. Going back and forth between $x^{+}({\boldsymbol{u}})$ and $x^{-}({\boldsymbol{u}})$ is called flipping the worm. Except for the tiles in the filled worm, all the tiles in $x^{+}({\boldsymbol{u}})$ and $x^{-}({\boldsymbol{u}})$ are the same. The following is the first part of the proof of Theorem 1.1.
Proposition 4.9.
If ${\mathcal{V}}$ is a direction perpendicular to a $5$th root of unity, then the Penrose tiling dynamical system $(X,\mathbb{R}^{2},T)$ is not expansive in the direction ${\mathcal{V}}$.
Proof.
Suppose ${\mathcal{V}}=\{{\boldsymbol{v}}_{j}\}^{\perp}$. Let $x^{+}\in\mathsf{X}_{2}$ and fix $r>0$. By applying a rotation and a translation we can assume that the single worm in $x^{+}$ is in a direction ${\boldsymbol{v}}_{j}^{\perp}$ that lies outside ${\mathcal{V}}^{r}$. Let $x^{-}$ be $x^{+}$ with the worm flipped. Then $x^{+}[{\mathcal{V}}^{r}]=x^{-}[{\mathcal{V}}^{r}]$ but $x^{+}\neq x^{-}$. It follows from Definition 3.12 and Proposition 3.13 that $(\mathsf{X},\mathbb{R}^{2},T)$ is not expansive in the direction ${\mathcal{V}}$. ∎
Figure 7. (a) A $5$-fold crossing in $y({\boldsymbol{u}})\in Y_{10}$ with some surrounding $3$-fold crossings, (b) a small perturbation $y({\boldsymbol{u}}^{\prime})\in Y_{1}$ of the $5$-fold crossing, and (c) the decagon tiling in rotation 00 together with worm filling directions: ,+,,,+-,+,-,-,+- , + , - , - , +.
Case B: Here, $y=y({\boldsymbol{u}})$ has a single $5$-fold crossing at ${\boldsymbol{b}}\in\mathbb{R}^{2}$. By translation, we can assume ${\boldsymbol{b}}={\boldsymbol{0}}$ so the spines are $\ell_{j,0}$, $j=0,\dots,4$. All the crossings along these spines, except at ${\boldsymbol{0}}$, are wide and narrow $3$-fold, and all the other crossings in $y({\boldsymbol{u}})$ are $2$-fold (see [6]). In $y({\boldsymbol{u}})^{*}$ the dual of the $5$-fold crossing is a regular decagon, and there are unfilled half-worms of hexagons radiating from its faces. This configuration is called an unfilled cartwheel, and we call a direction ${\boldsymbol{v}}_{j}^{\perp}$ line $l_{j}$ a spoke if it goes through the middle of a pair of half-worms.
Clearly $y({\boldsymbol{u}})^{*}$ is invariant under rotations around ${\boldsymbol{0}}$ by multiples of $102\pi/$ 10. We fill the unfilled cartwheel with $p_{5}$ tiles to obtain a Penrose tiling $x({\boldsymbol{u}})^{0}\in\mathsf{X}$ as follows. We fill the half-worm pairs, in the order ${\boldsymbol{v}}_{0}^{\perp},\dots,{\boldsymbol{v}}_{4}^{\perp}$, with a pattern of directions ,+,,,+-,+,-,-,+- , + , - , - , +. Each counter-clockwise rotation of $x({\boldsymbol{u}})^{0}$ by $102\pi/$ 10 acts on the sign pattern by flipping all the directions and cycling the pattern by moving the last direction to the front. So, for example, a single rotation of $x({\boldsymbol{u}})^{0}$, which we denote $x({\boldsymbol{u}})^{1}$, has the sign pattern ,+,,+,+-,+,-,+,+- , + , - , + , +. There are $10$ such patterns which yield the 10 distinct rotations of the decagon filling, and yield $10$ distinct Penrose tilings $x({\boldsymbol{u}})^{0},\dots,x_{9}({\boldsymbol{u}})\in\mathsf{X}$, obtained from filling the unfilled cartwheel in $y({\boldsymbol{u}})\in Y_{10}$. Off this unfilled cartwheel these are all the same. These $10$ Penrose tilings are the $10$ distinct $\varphi$ preimages of $y({\boldsymbol{u}})\in Y_{10}$, (see [6]).
5. Proof of Theorem 1.1
5.1. Sturmian sequences
Recall that $\alpha:=\gamma-1\approx 0.6180$. The slope $\alpha$ irrational rotation is the Kronecker system $(\mathbb{T},\mathbb{Z},R_{\alpha})$ defined by $R_{\alpha}u:=u+\alpha\!\!\mod 1$. It satisfies $\Sigma_{R_{\alpha}}=\mathbb{Z}[\alpha]/\mathbb{Z}\subseteq\mathbb{T}\sim\widehat{\mathbb{Z}}$, where $\mathbb{Z}[\alpha]$ is the ring generated by $\alpha$, and $\mathbb{Z}[\alpha]/\mathbb{Z}$ is its projection to the circle. A Sturmian sequence $z=(z_{n})_{n\in\mathbb{Z}}\in\{0,1\}^{\mathbb{Z}}$ is a coding of $u\in\mathbb{T}$ by $R_{\alpha}$. It is defined by $z_{n}=z(u)_{n}:=\chi_{I}(R^{n}_{\alpha}u)$, where $I$ is either $I^{+}:=[1-\alpha,1)$ or $I^{-}:=(1-\alpha,1]$. We let $Z=\{(z(u)_{n})_{n\in\mathbb{Z}}:u\in\mathbb{T}\}$, with $\phi:Z\to\mathbb{T}$ being the factor map satisfying $\phi(z(u))=u$. The following lemma summarizes some well known facts.
Lemma 5.1.
The $\mathbb{Z}$ Sturmian subshift $(Z,\mathbb{Z},S)$ (not SFT) is strictly ergodic almost automorphic with $\Sigma_{S}=\mathbb{Z}[\alpha]/\mathbb{Z}$, $\mathfrak{M}_{T}=\{1,2\}$, and $\mathbb{T}_{1}:=\mathbb{T}\backslash\{R^{n}_{\alpha}0:n\in\mathbb{Z}\}$.
Most of this can be found in [9], with explicit formula for $\phi(z)$ in [31]. We call the tensor square $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$ of $(Z,\mathbb{Z},S)$ the $\mathbb{Z}^{2}$ Sturmian subshift.
Lemma 5.2.
The $\mathbb{Z}^{2}$ Sturmian subshift $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$ is strictly ergodic almost automorphic with $\Sigma_{S\otimes S}=(Z[\alpha]/\mathbb{Z})\times(Z[\alpha]/\mathbb{Z})$, ((((i.e., an almost $1:1$ etension of $R_{\alpha}\otimes R_{\alpha}$)))), $\mathfrak{M}_{T}=\{1,2,4\}$, and $\mathbb{T}^{2}_{1}:=\mathbb{T}^{2}\backslash\{(R_{\alpha}\otimes R_{\alpha})^{\boldsymbol{n}}{\boldsymbol{0}}:{\boldsymbol{n}}\in\mathbb{Z}^{2}\}$. The factor map is $\phi\otimes\phi$.
5.2. Grid patches
Let
${{\boldsymbol{f}}}_{0}$ $:=(0,\sec(2\pi/20))=(0,1+\sqrt{5})\approx(0,3.236),{\rm\ and}$
${{\boldsymbol{f}}}_{1}$ $:=(1,-\tan(2\pi/20))=(1,\sqrt{5+2\sqrt{5}})\approx(-1,3.078),$ (5.1)
and let $A$ be the matrix with ${\boldsymbol{f}}_{0}$ and ${\boldsymbol{f}}_{1}$ as columns. We will define a rhombic tile $R$ by
$R:={\boldsymbol{f}}_{1}\wedge{\boldsymbol{f}}_{0}=\{t_{1}{{\boldsymbol{f}}}_{1}+t_{0}{{\boldsymbol{f}}}_{0}:t_{0},t_{1}\in[0,1]\}\subseteq{\mathbb{R}}^{2}=A[0,1]^{2}.$ (5.2)
[…]
$R_{\boldsymbol{n}}:=T^{-((-u_{0}+n_{0}){\boldsymbol{f}}_{0}+(-u_{1}+n_{1}){{\boldsymbol{f}}}_{1})}R=R+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime}),$ (5.3)
where in the second equality, we assume without loss of generality that ${\boldsymbol{u}}^{\prime}:=(u_{0},u_{1})\in[0,1)^{2}$.
Note that ${{\boldsymbol{0}}}\in R_{{\boldsymbol{0}}}$. We define the grid patch $g_{{\boldsymbol{n}}}({{\boldsymbol{u}}})$ to be the patch in $y({{\boldsymbol{u}}})$ supported on the rhombus $R_{{\boldsymbol{n}}}$ (see Figure 4(b)).
Definition 5.3.
Let $y({\boldsymbol{u}})\in Y_{1}$ be a nonsingular grid tiling (all crossings $2$-fold). For ${{\boldsymbol{n}}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, define the symbol of the grid patch $g_{{\boldsymbol{n}}}({{\boldsymbol{u}}})$, denoted by $\sigma(g_{{\boldsymbol{n}}}({{\boldsymbol{u}}}))=z_{n_{0}}\otimes z^{\prime}_{n_{1}}\in\{0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1\}$ as follows: If a single $4$-grid line passes through $R_{{\boldsymbol{n}}}$, put $z_{n_{0}}=0$, but if two $4$-grid lines pass through $R_{{\boldsymbol{n}}}$, put $z_{n_{0}}=1$. Similarly, if a single $2$-grid line passes through $R_{{\boldsymbol{n}}}$, put $z^{\prime}_{n_{1}}=0$, but if two $2$-grid lines pass through $R_{{\boldsymbol{n}}}$, put $z^{\prime}_{n_{1}}=1$.
Lemma 5.4.
Suppose that ${{\boldsymbol{u}}}\in\mathbb{T}^{5}_{0,1}$ satisfies $u_{0}=u_{1}=0$. Then $z_{0}\otimes z^{\prime}_{0}=\sigma(g_{{\boldsymbol{0}}}({{\boldsymbol{u}}}))$ is given by $z_{0}=\chi_{[1-\alpha,1)}(u_{4})$ and $z^{\prime}_{0}=\chi_{[1-\alpha,1)}(u_{2})$. Moreover, $\{\sigma(g_{{\boldsymbol{n}}}({{\boldsymbol{u}}}))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}=z(u_{4})\otimes z(u_{2})$.
Proof.
First note that $g_{{\boldsymbol{0}}}({{\boldsymbol{u}}})$ is supported on $R_{{\boldsymbol{0}}}$. A calculation (4.1) shows that $\ell_{2,0}(1-\alpha)$ passes through the upper right vertex of $R_{{\boldsymbol{0}}}$. If $z^{\prime}_{0}=0$, Definition 5.3 says that $\ell_{2,0}(u_{2})$ is the only $2$-grid line through $R_{{\boldsymbol{0}}}$. Since it enters $R_{{\boldsymbol{0}}}$ through the bottom edge (and exits through the top), it lies to the left of $\ell_{2,0}(1-\alpha)$. Thus $u_{2}\in[0,1-\alpha)$. Likewise, if $z^{\prime}_{0}=1$, there is a second 2-grid line $\ell_{2,-1}(u_{2})$ that enters $R_{{\boldsymbol{0}}}$ through the left edge and exits through the top. This means $\ell_{2,0}(u_{2})$ enters $R_{{\boldsymbol{0}}}$ to the right of $\ell_{2,0}(1-\alpha)$ (and exits through the right edge). In both cases, we have $z^{\prime}_{0}=\chi_{[1-\alpha,1)}(u_{2})$. The same argument shows $z_{0}=\chi_{[1-\alpha,1)}(u_{4})$. See Figure 8.
Figure 8. (a) In the case $z^{\prime}_{0}=0$, a single 2-grid line, $\ell_{2,0}(u_{2})$, crosses $R_{{\boldsymbol{0}}}$, so that $u_{2}\in[0,1-\alpha)$ since $\ell_{2,0}(u_{2})$ is left of $\ell_{2,0}(1-\alpha)$. (b) In the case $z^{\prime}_{0}=1$, two 2-grid lines cross $R_{{\boldsymbol{0}}}$ and $u_{2}\in[1-\alpha,1)$. (c) The singular case ($u_{2}=0$).
For the second part, $u_{0}=u_{1}=0$ implies $R_{\boldsymbol{n}}=T^{n_{0}{{\boldsymbol{f}}}_{1}+n_{1}{{\boldsymbol{f}}}_{0}}R$, so $z_{n_{0}}\otimes z^{\prime}_{n_{1}}=\sigma(g_{{\boldsymbol{0}}}(K^{-{\boldsymbol{s}}_{\boldsymbol{n}}}{{\boldsymbol{u}}}))=\sigma(g_{{\boldsymbol{0}}}(\{-W{\boldsymbol{s}}_{\boldsymbol{n}}+{\boldsymbol{u}}\})$. By a computation, $W{{\boldsymbol{f}}}_{0}=(0,1,\alpha,-\alpha,-1)$ and $W{{\boldsymbol{f}}}_{1}=(1,0,-1,-\alpha,\alpha)$, and $-W{\boldsymbol{s}}_{\boldsymbol{n}}+{\boldsymbol{u}}=(n_{0},n_{1},-n_{0}+n_{1}\alpha+u_{2},-(n_{0}+n_{1})\alpha+u_{3},n_{0}\alpha-n_{1}+u_{4})$, since ${\boldsymbol{u}}=(0,0,u_{2},u_{3},u_{4})$. By the definition of a Sturmian sequence, $z^{\prime}_{n_{1}}=\chi_{[1-\alpha,1)}(\{n_{1}\alpha+u_{2}\})=\chi_{[1-\alpha,1)}(R_{\alpha}^{n_{1}}u_{2})={z}(u_{2})_{n_{1}}$. The same argument shows $z_{n_{0}}={z}(u_{4})_{n_{0}}$. ∎
5.3. Duals of grid patches
Lemma 5.5.
Let $x\in\mathsf{X}_{1}$ be a nonsingular Penrose tiling with $x=y({\boldsymbol{u}})^{*}$ for $y=y({\boldsymbol{u}})\in Y_{1}$. Suppose that ${\boldsymbol{b}}_{\boldsymbol{n}}\in v(y)$ satisfies ${\boldsymbol{b}}_{\boldsymbol{n}}=\ell_{0,n_{0}}\cap\ell_{1,n_{1}}$, ${\boldsymbol{n}}=(n_{0},n_{1})$, so ${\boldsymbol{b}}_{\boldsymbol{n}}$ is the lower left vertex of the rhombus $R_{\boldsymbol{n}}$ in (5.3). Let $E_{\boldsymbol{n}}\in y$ be the tile in $g_{\boldsymbol{n}}({\boldsymbol{u}})$ with vertex ${\boldsymbol{b}}_{\boldsymbol{n}}$. The dual $D_{\boldsymbol{n}}={\boldsymbol{b}}_{\boldsymbol{n}}^{*}\in x$ of ${\boldsymbol{b}}_{\boldsymbol{n}}$ is a type ${\boldsymbol{v}}^{\prime}_{0}\wedge{\boldsymbol{v}}^{\prime}_{1}$ tile from $p_{5}$ in $x$, and the dual ${\boldsymbol{d}}_{\boldsymbol{n}}=E_{\boldsymbol{n}}^{*}$ of the tile $E_{\boldsymbol{n}}\in y$ is the upper right vertex of $D_{\boldsymbol{n}}$. Then $||{\boldsymbol{b}}_{\boldsymbol{n}}-{\boldsymbol{d}}_{\boldsymbol{n}}||<(1/5)(1+\sqrt{5})\approx 0.647$.
Proof.
Use Lemma 4.6 to translate ${\boldsymbol{b}}_{\boldsymbol{n}}$ to the origin, and making it ${\boldsymbol{b}}_{\bf 0}$, by replacing ${\boldsymbol{u}}$ with $K^{-{\boldsymbol{b}}_{\boldsymbol{n}}}{\boldsymbol{u}}$. Then $\{{\boldsymbol{u}}\}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$ with $u_{2},u_{4}\in[0,1)$. We show $||{\boldsymbol{d}}_{\bf 0}||<(1/2)(1+\sqrt{5})$. By (4.3), ${\boldsymbol{d}}_{\bf 0}=E^{*}_{\bf 0}:=(2/5)W^{t}({\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})+\{{\boldsymbol{u}}\})$ for any ${\boldsymbol{s}}\in{\rm int}(E_{\bf 0})$. Making ${\boldsymbol{s}}>0$ small, we see from (4.2) that ${\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})={{\boldsymbol{0}}}$ for ${\boldsymbol{s}}\in{\rm int}(E_{\bf 0})$, since ${\boldsymbol{m}}$ is constant on tiles. Thus $||{\boldsymbol{d}}_{\bf 0}||=(2/5)||W^{t}\{{\boldsymbol{u}}\}||\leq(2/5)||{\boldsymbol{v}}_{2}+{\boldsymbol{v}}_{3}+{\boldsymbol{v}}_{4}||=(2/5)(1+\sqrt{5})/2=(1/5)(1+\sqrt{5})$. ∎
Let $x\in\mathsf{X}$ (or more generally, $x\in X_{p_{5}}$). A ${\boldsymbol{v}}_{j}^{\perp}$-trail in $x$ is an infinite sequence of tiles $\{\dots,D_{-1},D_{0},D_{1},\dots\}$ in $x$ such that each $D_{i+1}$ is adjacent across to $D_{i}$ across a type ${\boldsymbol{v}}^{\prime}_{j}$ edge in the direction $R_{2\pi/4}{\boldsymbol{v}}_{j}$. For $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$ every ${\boldsymbol{v}}_{j}^{\perp}$-trail is the dual of the sequence of $2$-fold crossings along a grid line $\ell_{j,k}(u_{k})$ in $y({\boldsymbol{u}})$ for some $k\in\mathbb{Z}$. A ${\boldsymbol{v}}_{j}^{\perp}$-trail segment is a contiguous finite subsequence $\{D_{k},D_{k+1},\dots,D_{k+n-1}\}$ of a ${\boldsymbol{v}}_{j}^{\perp}$-trail.
Lemma 5.6.
Fix a Penrose tiling $x\in\mathsf{X}$ and let $x^{\prime}$ be the set of all type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}_{1}^{\prime}$ tiles in $D\in x$. Then there is a bijective map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}:\mathbb{Z}^{2}\to x^{\prime}$ such that for each ${\boldsymbol{n}}\in\mathbb{Z}^{2}$, a ${\boldsymbol{v}}_{0}^{\perp}$-trail segment connects $D_{\boldsymbol{n}}$ to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$ and a ${\boldsymbol{v}}_{1}^{\perp}$-trail segment connects $D_{\boldsymbol{n}}$ to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$. The map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}$ is unique for $x\in\mathsf{X}_{1}$ if we assume $D_{\bf 0}={\boldsymbol{b}}_{\bf 0}^{*}$, or for $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ if we assume $D_{\bf 0}=({\boldsymbol{b}}^{\prime}_{\bf 0})^{*}$ with $x^{\prime}\in\mathsf{X}_{1}$ and $\rho(x,x^{\prime})$ sufficiently small.
Proof.
If $x\in\mathsf{X}_{1}$ is nonsingular, the map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}$ defined in Lemma 5.5 satisfies the lemma. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ is singular, there is a sequence $x_{k}\in\mathsf{X}_{1}$ so that $x_{k}\to x$ in the metric $\rho$, (i.e., $\mathsf{X}=\overline{\mathsf{X}_{1}}$, Lemma 4.8). Let $x_{k}=y({\boldsymbol{u}}_{k})^{*}$ and let ${\boldsymbol{n}}\mapsto D^{k}_{\boldsymbol{n}}$ be the map defined in Lemma 5.5. It satisfies $D^{k}_{\boldsymbol{n}}={\boldsymbol{b}}^{k}_{\boldsymbol{n}}$ for ${\boldsymbol{b}}^{k}_{\boldsymbol{n}}\in v(y({\boldsymbol{u}}_{k}))$. Let $r_{n}>0$ be an increasing sequence, and for each $n$, choose $K_{n}$ large enough that $k\geq K_{n}$ implies $x_{k}[B_{r_{n}}({\bf 0})]=x[B_{r_{n}}({\bf 0})]$. Define $D_{\boldsymbol{n}}:=D^{k}_{\boldsymbol{n}}$ where $r_{n}>||{\boldsymbol{n}}||$ and $k\geq K_{n}$. ∎
Definition 5.7.
For a Penrose tiling $x\in\mathsf{X}$ and ${\boldsymbol{n}}\in\mathbb{Z}^{2}$, we define a patch $w_{\boldsymbol{n}}(x)$ in $x$, called a Wang patch, as follows: Starting at $D_{\boldsymbol{n}}$, go forward along the ${\boldsymbol{v}}^{\perp}_{0}$-trail segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$ (Lemma 5.6), then go forward along the ${\boldsymbol{v}}^{\perp}_{1}$-trail segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$, then go backward along the ${\boldsymbol{v}}^{\perp}_{0}$-trail segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$, and finally go backward along the ${\boldsymbol{v}}^{\perp}_{1}$-trail segment back to $D_{\boldsymbol{n}}$. The patch $w_{\boldsymbol{n}}(x)$ is defined to be this loop of closed trail segments, together with all the tiles inside it.
The Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ in $x\in\mathsf{X}$ cover $\mathbb{R}^{2}$ in the sense that $x=\bigcup_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}w_{\boldsymbol{n}}(x)$. We call the four type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}_{1}^{\prime}$ tiles in $w_{\boldsymbol{n}}(x)$ its corners. The following is a consequence of Lemma 5.6.
Lemma 5.8.
For $x\in\mathsf{X}_{1}$, let $x=y({\boldsymbol{u}})^{*}$ where $y({\boldsymbol{u}})\in Y_{1}$. Then each Wang patch in $x$ is the dual of the corresponding grid patch: $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$.
Figure 9. The $24$ Wang patches.
Proposition 5.9 (Raphael Robinson, see Exercise 11.1.2 in [6]).
Up to translation, there are $24$ Wang (proto-) patches among all the Penrose tilings $x\in\mathsf{X}$.
We denote these Wang (proto-) patches by $\{w^{a}:a\in{\mathcal{A}}\}$ where ${\mathcal{A}}=\{0,\dots,23\}$. These $24$ Wang patches are shown in Figure 9. Pictures of these patches appeared in the first author’s dissertation [17], and also in [16].
Figure 10. The bifurcation diagram, showing the parameter values $(u_{2},u_{4})\in\mathbb{T}^{2}\sim[0,1]^{2}$ corresponding the 24 Wang patches. Each cell is labeled with $a\in{\mathcal{A}}$ for the corresponding Wang patch $w^{a}$ (Figure 9). The vertices along each edges of the square are at 00, $2-\gamma\approx 0.381967$, $\gamma-1\approx 0.618034$, $4-2\gamma\approx 0.763932$, and $1$. The coarser division of $[0,1]^{2}$ into four cells by the lines $u_{2}=2-\gamma$ and $u_{4}=2-\gamma$ corresponds to the values $0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1$ of the symbol $\sigma(y({\boldsymbol{u}}))$ where ${\boldsymbol{u}}=(0,0,u_{2},u_{3},u_{4})$ with $u_{2}+u_{3}+u_{4}=0$, corresponding to the $\mathbb{Z}^{2}$ Sturmian shift $S\otimes S$.
Proof of Proposition 5.9.
By Lemma 5.8, all the Wang patches $w^{a}$, up to translation, can be obtained as duals $g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$ of all the grid patches $g_{\boldsymbol{n}}({\boldsymbol{u}})$ that occur in all $y({\boldsymbol{u}})\in Y_{1}$. This means all crossings in $y({\boldsymbol{u}})$ are 2-fold, and in particular, this holds for any grid patch in $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})$. We will call a grid patch that has only 2-fold crossings a nonsingular grid patch. Now, if $g_{\boldsymbol{n}}({\boldsymbol{u}})$ is a nonsingular grid patch in $y({\boldsymbol{u}})\in Y$, (i.e., $y({\boldsymbol{u}})$ is not necessarily nonsingular), then since $Y_{1}$ is dense $G_{\delta}$, there exists a small perturbation of ${\boldsymbol{u}}$, which does not change $g_{\boldsymbol{n}}({\boldsymbol{u}})$ enough to change its dual $g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$, so that with this new ${\boldsymbol{u}}$, we have $y({\boldsymbol{u}})\in Y_{1}$.
Given any grid patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})\in Y$ we can translate it to the origin, using Lemma 4.6 to adjust ${\boldsymbol{u}}$, so that for this new ${\boldsymbol{u}}$, the patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ is translated to the origin and becomes $g_{\boldsymbol{0}}({\boldsymbol{u}})$, with its support satisfying $R_{\boldsymbol{0}}=R$, where $R$ is as in (5.2). The new value of ${\boldsymbol{u}}$ thus has the form ${\boldsymbol{u}}=(0,0,u_{2},u_{3},u_{4})$, where we assume without loss of generality that $(u_{2},u_{3})\in[0,1]^{2}$ with $u_{2}+u_{3}+u_{4}=0$. Thus we think of $g_{\boldsymbol{0}}({\boldsymbol{u}})$ as depending on $(u_{2},u_{4})\in[0,1]^{2}$, but we also observe that $g_{\boldsymbol{0}}({\boldsymbol{u}})$ is the same for $(u_{2},0)$ and $(u_{2},1)$, and for $(0,u_{4})$ and $(1,u_{4})$, so we can also regard $(u_{2},u_{4})\in\mathbb{T}^{2}$.
Figure 10 shows a partition of $[0,1]^{2}$ into $24$ cells, $46$ edges, and $23$ vertices. We call this the bifurcation diagram for Wang patches, noting that the same diagram, called a Markov partition for Wang tiles, was obtained in [16]. The cells in the bifurcation diagram are defined by the condition that a point $(u_{2},u_{4})$ is in the interior of a cell if and only if the corresponding grid patch $g_{{\boldsymbol{0}}}({\boldsymbol{u}})$ is nonsingular. It follows that the dual $g_{{\boldsymbol{0}}}({\boldsymbol{u}})^{*}$ is a constant Wang patch $w^{a}$ for all $(u_{2},u_{4})$ in the interior of each cell. By the discussion in the previous two paragraphs, the duals $g_{{\boldsymbol{0}}}({\boldsymbol{u}})^{*}$ range over all possible Wang patches $w^{a}$ up to translation. We can see each of these Wang patches by drawing the dual for the grid patch $g_{\boldsymbol{0}}({\boldsymbol{u}}_{0})$ corresponding to a single point ${\boldsymbol{u}}_{0}$ in each cell. We find that each cell results in a distinct Wang patch. Thus there are 24 distinct Wang patches, which are the Wang patches $w^{a}$, $a\in{\mathcal{A}}$, shown in Figure 9. The numbering $a\in{\mathcal{A}}$ corresponds to the labels of the cells in Figure 10. In particular, these are the 24 Wang patches described in [11].
It remains to show that the bifurcation diagram is correct. Figure 10 shows five vertices along each side of $[0,1]^{2}$ located at $\eta_{0}=0$, $\eta_{1}:=2-\gamma\approx 0.381967$, $\eta_{2}:=\gamma-1\approx 0.618034$, $\eta_{3}:=4-2\gamma=3-\sqrt{5}\approx 0.763932$, and $\eta_{4}=1$. We will now locate all the possible multiple crossings in $g_{\boldsymbol{0}}({\boldsymbol{u}})$. The remainder of the proof is organized in five steps.
Step 1. It is easy to see that there are $5$-fold crossings $g_{\boldsymbol{0}}({\boldsymbol{u}})$ exactly when $(u_{2},u_{4})$ in $[0,1]^{2}$ is one of the nine vertices corresponding to the partition of $[0,1]^{2}$ into four cells by the vertical line $u_{2}=\eta_{1}$ and the horizontal line $u_{4}=\eta_{1}$, (in $\mathbb{T}^{2}$ these identify to four vertices). In subsequent discussions we implicitly ignore these nine points, and look only for $3$-fold crossings in $g_{\boldsymbol{0}}({\boldsymbol{u}})$.
Altogether, there are ten possible types of $3$-fold crossings, which we denote $012$, (meaning a crossing of a 00-, $1$- and a $2$-grid line), $013$, $014$, $023$, $024$, $034$ $123$, $124$, $134$, and $234$.
Step 2. It is easy to see that if $(u_{2},u_{4})$ lies along the left or right side of $[0,1]^{2}$ (on the line $u_{2}=0$ or $u_{2}=1$) then there are $2$-grid lines through the lower left and lower right corners of $R_{\boldsymbol{0}}$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$, as shown in Figure 8(c). Similarly, for points along the vertical line $u_{2}=\eta_{1}$, there are $2$-grid lines through the upper left and upper right corners of $R_{\boldsymbol{0}}$. These account for all possible $3$-fold crossings type of $012$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$. In a similar way, for $(u_{2},u_{4})$ along any of the three horizontal lines $u_{2}=0$, $u_{2}=\eta_{1}$ or $u_{2}=1$, there are $3$-fold crossings of type $014$ at corners of $R_{\boldsymbol{0}}$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$, and there can be no other crossings of this type.
Step 3. For all $(u_{2},u_{4})$ along the five lines of slope $-1$ in the bifurcation diagram, (i.e., the lines $u_{2}+u_{4}=\eta_{1}$, $u_{2}+u_{4}=\eta_{2}$, $u_{2}+u_{4}=1$, $u_{2}+u_{4}=2\eta_{2}$ and $u_{2}+u_{4}=2\eta_{3}$) there are $3$-fold crossings type $013$ at corners of $R_{\boldsymbol{0}}$. For two of these, $u_{2}+u_{4}=\eta_{1}$ and $u_{2}+u_{4}=2\eta_{2}$, the $3$-grid lines cross the upper left and lower right corners. For the other three, the $3$-grid lines cross the lower left and upper right corners of $R_{\boldsymbol{0}}$. No other crossings of type $013$ are possible.
Step 4. We show that for $(u_{2},u_{4})\in[0,1]^{2}$, all possible type $234$ crossings occur along (two of the) slope $-1$ lines from Step 3: $u_{2}+u_{4}=\eta_{1}$ and $u_{2}+u_{4}=2\eta_{2}$. By Section 4.5, Case A, all $3$-fold crossings in $y({\boldsymbol{u}})$ occur along spines with both wide and narrow $3$-fold crossings (Figure 6(a)). In Step 3 above, the spine is a $3$-grid line. The type $013$ in Step 3 are wide crossings, so the narrow crossings are type $234$. These occur in $g_{\boldsymbol{0}}({\boldsymbol{u}})$ in the cases where the $3$-grid line goes through the lower left and upper right corners of $R_{\boldsymbol{0}}$: $u_{2}+u_{4}=\eta_{1}$ and $u_{2}+u_{4}=2\eta_{2}$.
Step 5. Finally, we discuss the six types of $3$-fold crossings that have the form $0ij$ or $1ij$ for $i,j=2,3,4$, $i<j$. Any such crossing occurs along an edge of $R_{\boldsymbol{0}}$. By (4.1), a grid line $\ell_{i,k}(u_{i})$ has the equation ${\boldsymbol{v}}_{i}\cdot(x,y)=(-u_{i}+k)$. So if $\ell_{i,k}(u_{i})\cap\ell_{j,k^{\prime}}(u_{j})=\{(x,y)\}$ then $V_{i,j}(x,y)=(-u_{i}+k,-u_{j}+k^{\prime})$, where $V_{i,j}$ is the $2\times 2$ matrix with rows ${\boldsymbol{v}}_{i}$ and ${\boldsymbol{v}}_{j}$. Recalling that we identify $(x,y)$ with $[x,y]^{t}$, we define an affine map $A_{i,j,k,k^{\prime}}(x,y):=M_{i,j}V_{i,j}(x,y)-(k,k^{\prime})$, where
$M_{2,4}:=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right),\ \ \ M_{2,3}:=\left(\begin{matrix}1&0\\ -1&-1\end{matrix}\right),\ \ \ M_{3,4}:=\left(\begin{matrix}-1&-1\\ 0&1\end{matrix}\right).$ (5.4)
Then $A_{i,j,k,k^{\prime}}$ maps a crossing $(x,y)$ of $\ell_{i,k}(u_{i})$ and $\ell_{j,k^{\prime}}(u_{j})$ ($i,j=2,3,4$, $i<j$) to a point $(u_{2},u_{4})$ in the plane of the bifurcation diagram. The matrices $M_{i,j}$ correspond to the fact that $u_{2}+u_{3}+u_{4}=0$. When the map $A_{i,j,k,k^{\prime}}$ is applied to the boundary $\partial R_{\boldsymbol{0}}$ of $R_{\boldsymbol{0}}$ the image is a parallelogram in the $(u_{2},u_{4})$-plane whose intersections with $[0,1]^{2}$ are the values of $(u_{2},u_{4})$ in the bifurcation diagram that correspond to $3$-fold crossings of type $0ij$ or $1ij$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$.
For each type $ij\in\{24,23,34\}$, Figure 11 shows $A_{i,j,k,k^{\prime}}(\partial R_{\boldsymbol{0}})$ for each of the values of $(k,k^{\prime})$ so that $A_{i,j,k,k^{\prime}}(\partial R_{\boldsymbol{0}})\cap[0,1]^{2}\neq\emptyset$. The caption describes which sides correspond to type $0ij$ and $1ij$ and which $(k,k^{\prime})$ are used.
We conclude by observing that steps 1 - 5 account for all possible $3$-fold and $5$-fold crossings in $g_{\boldsymbol{0}}({\boldsymbol{u}})$, and also account for all the segments in Figure 10.. This shows that the bifurcation diagram is correct. ∎
Figure 11. Images of $\partial R_{\boldsymbol{0}}$ under $A_{i,j,k,k^{\prime}}$ in the $(u_{2},u_{4})$-plane are parallelograms (black). Only those choices of $k,k^{\prime}$ resulting in images that meet $[0,1]^{2}$ (grey) are shown. (a) Images under $A_{2,4,0,0}$, $A_{2,4,0,1}$, and $A_{2,4,1,0}$. The slope $-\gamma$ sides correspond to type $024$ and slope $-1/\gamma$ sides to type $124$. (b) Images under $A_{2,3,0,-1}$, $A_{2,3,-1,-1}$, and $A_{2,3,0,-2}$. The slope $-\gamma$ sides correspond to type $123$, and slope 00 sides (including part of $\partial[0,1]^{2}$) correspond to type $023$. (c) Images under $A_{3,4,-2,0}$, $A_{3,4,-1,0}$, and $A_{3,4,-1,-1}$. The slope $-1/\gamma$ sides correspond to type $034$ and vertical sides (including part of $\partial[0,1]^{2}$) correspond to type $134$.
Definition 5.10.
Fix $x\in\mathsf{X}$. For each $n_{1}\in\mathbb{Z}$ we define the $n_{1}$th bent 00-grid line in $x$, denoted $\lambda_{0,n_{1}}$, to be the piecewise linear curve connecting the points $\{{\boldsymbol{d}}_{(n_{0},n_{1})}\}_{n_{0}\in\mathbb{Z}}$ (where the points ${\boldsymbol{d}}_{\boldsymbol{n}}$ are as defined in Lemma 5.5). Similarly, for each $n_{0}\in\mathbb{Z}$, we define the $n_{0}$th bent $1$-grid line in $x$, denoted $\lambda_{1,n_{0}}$, to be the piecewise linear curve connecting the vertices $\{{\boldsymbol{d}}_{(n_{0},n_{1})}\}_{n_{1}\in\mathbb{Z}}$.
The next lemma shows that bent 00- and $1$-grid lines $\lambda_{1,n_{0}}$ and $\lambda_{n_{1},0}$ closely follow the corresponding grid lines $\ell_{1,n_{0}}$ and $\ell_{n_{1},0}$.
Lemma 5.11.
Suppose $x\in\mathsf{X}_{1}$ and $x=y({\boldsymbol{u}})^{*}$ for $y({\boldsymbol{u}})\in Y_{1}$. Then $||\ell_{0,n_{1}}-\lambda_{0,n_{1}}||,||\ell_{1,n_{0}}-\lambda_{1,n_{0}}||<2/3$ (in the sense of the Euclidean uniform closeness of the two curves). The same results hold for singular Penrose tilings $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, where $y({\boldsymbol{u}})=\varphi(x)$.
Proof.
For $x\in\mathsf{X}_{1}$, this follows from Lemma 5.5. For $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, we let $x_{k}=y({\boldsymbol{u}}_{k})^{*}\to x$, where $y({\boldsymbol{u}}_{k})\in Y_{1}$ (see Lemma 4.8) can be chosen so that $\ell_{0,n_{1}}(u_{0})=\ell_{0,n_{1}}(u^{k}_{0})$ and $\ell_{0,n_{1}}(u_{0})=\ell_{0,n_{1}}(u^{k}_{0})$ for all $||{\boldsymbol{n}}||$ small. ∎
The bent 00- and $1$-grid lines $\{\lambda_{1,n_{0}}(x)\}_{n_{0}\in\mathbb{Z}}$ and $\{\lambda_{0,n_{1}}(x)\}_{n_{1}\in\mathbb{Z}}$ partition $\mathbb{R}^{2}$ into a tiling by tetragons (i.e., quadrilaterals):
Definition 5.12.
For $x\in\mathsf{X}$ and ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, let $U_{\boldsymbol{n}}(x)$ be the tile bounded by the tetragon with vertices ${\boldsymbol{d}}_{\boldsymbol{n}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$, and ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$, (the edges of $U_{\boldsymbol{n}}(x)$ are segments of the edge curves $\lambda_{1,n_{0}}$, $\lambda_{1,n_{0}+1}$, $\lambda_{0,n_{1}}$ and $\lambda_{0,n_{1}+1}$ in $x$). We call $U_{\boldsymbol{n}}(x)$ an unmarked tetragon Penrose tile, and call $u(x):=\{U_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ an unmarked tetragon Penrose tiling.
There is a unique tetragon tile $U_{\boldsymbol{n}}(x)$ in $u(x)$ corresponding to each Wang patch $w_{\boldsymbol{n}}(x)$ in $x$ (see Figure 12). In Section 6 we will show that there are, up to translation, $11$ different tetragon prototiles $\{U^{b}:b=0,\dots,10\}$, so some of the tetragon tiles correspond to more than one Wang patche $w^{a}$. We will return to these tilings in Section 6.
Figure 12. Swatch of Penrose tiling $x\in\mathsf{X}$ (a) with type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}^{\prime}_{1}$ tiles $D_{\boldsymbol{n}}$ shaded to show corners of Wang patches $w^{\boldsymbol{n}}(x)$. (b) Edge curves connect the upper right vertices ${\boldsymbol{d}}_{\boldsymbol{n}}$ of $D_{\boldsymbol{n}}$ delineating unmarked tetragon Wang tiles $U_{\boldsymbol{n}}$.
Definition 5.13.
For each Wang patch $w^{a}$, $a\in{\mathcal{A}}$, we define its symbol $\sigma(w^{j})=z_{0}\otimes z^{\prime}_{0}\in\{0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1\}$ as follows: If a single direction ${\boldsymbol{v}}_{2}^{\perp}$ trail segment passes through $w^{j}$, we put $z^{\prime}_{0}=0$, but if two distinct direction ${\boldsymbol{v}}_{2}^{\perp}$ trail segments pass through $w^{j}$, we put $z^{\prime}_{0}=1$. Similarly, if a single direction ${\boldsymbol{v}}_{4}^{\perp}$ trail segment passes through $w^{j}$, we put $z_{0}=0$, but if two distinct direction ${\boldsymbol{v}}_{4}^{\perp}$ trail segments pass through $w^{j}$, we put $z_{0}=1$, (compare Definition 5.3).
Lemma 5.14.
Let $x\in\mathsf{X}$ be a Penrose tiling covered by Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$. Then $z\otimes z^{\prime}=\{\sigma(w_{{\boldsymbol{n}}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence. Let $u_{2}=\phi(z)$, $u_{4}=\phi(z^{\prime})$, and let ${\boldsymbol{u}}_{0}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$. If $x\in\mathsf{X}_{1}$ then $y({\boldsymbol{u}}_{0})\in Y_{1}$ and $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$ for some unique ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ then $y({\boldsymbol{u}}_{0})^{*}$ has an unfilled worm or unfilled cartwheel, and there is a unique ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$ so that $x$ is $T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}})^{*}$ with the worm or cartwheel filled.
Proof.
First let $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$, so that $y({\boldsymbol{u}})\in Y_{1}$. By Definitions 5.3 and 5.13, $\sigma(g_{{\boldsymbol{n}}}({\boldsymbol{u}})^{*})=\sigma(g_{{\boldsymbol{n}}}({\boldsymbol{u}}))$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Thus $\{\sigma(w_{\boldsymbol{n}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence since $\{\sigma(g_{\boldsymbol{n}}({\boldsymbol{u}}))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is $\mathbb{Z}^{2}$ Sturmian by Lemma 5.4. For ${\boldsymbol{u}}_{0}$ as described in the lemma, $y({\boldsymbol{u}}_{0})$ and $y({\boldsymbol{u}})$ are translates since they read the same Sturmian sequence. Thus there is a unique ${\boldsymbol{t}}_{0}$ so that $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})$. For $y({\boldsymbol{u}})$ we have ${\boldsymbol{0}}\in R_{\boldsymbol{0}}$, whereas for $y({\boldsymbol{u}})^{*}$ the lower left vertex of $R_{\boldsymbol{0}}$ is ${\boldsymbol{0}}$. Thus $||{\boldsymbol{t}}_{0}||\leq||{\boldsymbol{f}}_{0}+{\boldsymbol{f}}_{1}||<4/3$, and (4.4) implies $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$.
If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ let $x_{n}\to x$ in $\mathsf{X}$. Let $(z\otimes z^{\prime})_{n}$ be the $\mathbb{Z}^{2}$ Sturmian sequence for $x_{n}$, and let ${\boldsymbol{u}}_{0,n}$ be the value of ${\boldsymbol{u}}_{0}$ (Lemma 5.4) for $(z\otimes z^{\prime})_{n}$. Then the definition of the tiling metric $\rho$ provides that $x_{n}\to x$ implies $(z\otimes z^{\prime})_{n}$ converges to a $\mathbb{Z}^{2}$ Sturmian sequence $z\otimes z^{\prime}$, which is the sequence read from $x$, (i.e., $z_{n_{0}}\otimes z^{\prime}_{n_{1}}=\sigma(w_{(n_{0},n_{1})}(x))$. If ${\boldsymbol{u}}=y^{-1}(\varphi(x))$ then since $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, $y({\boldsymbol{u}})^{*}$ has either an unfilled worm or an unfilled cartwheel. But also $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0,n}}y({\boldsymbol{u}}_{0,n})$ for some ${\boldsymbol{t}}_{0,n}$ with $||{\boldsymbol{t}}_{0,n}||<4/3$, and in fact ${\boldsymbol{t}}_{0,n}={\boldsymbol{t}}_{0}$ for all sufficently large $n$. So $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})$, but also $T^{{\boldsymbol{t}}_{n}}y({\boldsymbol{u}}_{0,n})^{*}=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0,n})^{*}\to x$. Thus $x$ is obtained from $y({\boldsymbol{u}})^{*}$ by filling in a worm or cartwheel. ∎
5.4. The proof of Theorem 1.1
The next two propositions, together with Proposition 4.9, when used in conjunction with Proposition 3.13, complete the proof of Theorem 1.1.
Proposition 5.15.
Let $x\in\mathsf{X}_{1}$ be a nonsingular Penrose tiling and let ${\mathcal{V}}\subseteq\mathbb{R}^{2}$ be a direction in $\mathbb{R}^{2}$ not perpendicular to either ${\boldsymbol{v}}_{0}$ or ${\boldsymbol{v}}_{1}$. Then there is an $r_{1}>0$ (independent of $x$ and ${\mathcal{V}}$) so that if $x[{\mathcal{V}}^{r_{1}}]=x^{\prime}[{\mathcal{V}}^{r_{1}}]$ for any (not necessarily nonsingular) Penrose tiling $x^{\prime}\in\mathsf{X}$, then $x=x^{\prime}$.
Proposition 5.16.
Let $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ be a singular Penrose tiling and let ${\mathcal{V}}\subseteq\mathbb{R}^{2}$ be a direction in $\mathbb{R}^{2}$ not perpendicular to any 5th root of unity. Then there is an $r_{2}>0$ (independent of $x$ and ${\mathcal{V}}$) so that if $x[{\mathcal{V}}^{r_{2}}]=x^{\prime}[{\mathcal{V}}^{r_{2}}]$ for any Penrose tiling $x^{\prime}\in\mathsf{X}$, then $x=x^{\prime}$.
The proofs of these Propositions are similar, but we separate them for conceptual clarity. We begin with the following two lemmas.
Lemma 5.17.
There exists an $r_{0}>0$ so that the following holds. Suppose $x\in\mathsf{X}$ with $x=T^{{\boldsymbol{t}}_{0}}x^{\prime}$ for $||{\boldsymbol{t}}_{0}||<4/3$. If $x[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]$ for $||{\boldsymbol{t}}^{\prime}_{0}||<4/3$ then ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$.
This is essentially the locally free property for $\mathsf{X}$, but on a new scale.
Proof.
For the unmarked Penrose prototiles $p_{5}$ we compute $\Delta_{p_{5}}=2/5$, (see Definition 3.2). For any prototile set $p$, $\Delta_{\gamma p}=\gamma{\Delta_{p}}$, so $\Delta_{\gamma^{3}p_{5}}>4/3$. Consider the MLD topological conjugacy $(C^{-1})^{3}:\mathsf{X}\to\gamma^{3}\mathsf{X}$ where $\gamma^{3}\mathsf{X}\subseteq X_{\gamma^{3}p_{5}}$. Let $r_{0}>0$ be the radius for the local map $(C^{-1})^{3}$, (see Proposition 3.5). Now assume $(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]=x[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]$. Let $x_{3}:=(C^{-1})^{3}x\in\gamma^{3}\mathsf{X}$. By conjugacy, $(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime}_{3}))[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}_{0}}x^{\prime}_{3}))[B_{r_{0}}({\boldsymbol{0}})]$. Since $||{\boldsymbol{t}}_{0}||,||{\boldsymbol{t}}^{\prime}_{0}||<4/3<\Delta_{\gamma^{3}p}$, Lemma 3.3(1) shows ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$. ∎
Lemma 5.18.
There is an $r^{\prime}_{1}>0$ so that for all $x\in\mathsf{X}_{1}$:
$\forall n_{0}\in\mathbb{Z}\text{\rm\ and }{\boldsymbol{s}}_{0}\in\lambda_{1,n_{0}},\exists n^{\prime}_{1}\in\mathbb{Z}\text{\rm\ so\ that }w_{(n_{0},n_{1})}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{0}),$ (5.5)
and
$\forall n_{1}\in\mathbb{Z}\text{\rm\ and }{\boldsymbol{s}}\in\lambda_{0,n_{1}},\exists n^{\prime}_{0}\in\mathbb{Z}\text{\rm\ so\ that }w_{(n_{0},n_{1})}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{)}.$ (5.6)
Proof.
There is an $r^{\prime}_{1}>0$ so that for any ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$ the following holds: If ${\boldsymbol{s}}_{0}\in\lambda^{\prime}_{0,n_{1}}=\overline{{\boldsymbol{d}}_{\boldsymbol{n}}\,{\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}}$ and ${\boldsymbol{s}}_{1}\in\lambda^{\prime}_{1,n_{0}}=\overline{{\boldsymbol{d}}_{\boldsymbol{n}}\,{\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}}$, then $w_{\boldsymbol{n}}(x)\cup w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{0})$ and $w_{\boldsymbol{n}}(x)\cup w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{1})$. Thus for any $n_{0}\in\mathbb{Z}$ and ${\boldsymbol{s}}\in\lambda_{1,n_{0}}(x)$, $B_{r^{\prime}_{1}}({\boldsymbol{s}})$ contains at least one Wang patch from the row $\{w_{(n_{0},n_{1})}(x)\}_{n_{1}\in\mathbb{Z}}$. This proves (5.5), and the proof of (5.6) is the same. ∎
Proof of Proposition 5.15.
For each ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, let $z_{n_{0}}\otimes z^{\prime}_{n_{1}}:=\sigma(w_{\boldsymbol{n}}(x))$. By the assumption that $x\in\mathsf{X}_{1}$, Lemma 5.14 implies that $z\otimes z^{\prime}\in Z\otimes Z$ is a $\mathbb{Z}^{2}$ Sturmian sequence, Moreover, Lemma 5.14 implies that if $u_{2}=\phi(z)$, $u_{4}=\phi(z^{\prime})$ and $u_{3}=\{-(u_{2}+u_{4})\}$, then for ${\boldsymbol{u}}_{0}=(0,0,u_{2},u_{3},u_{4})$ we have $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$ for $y({\boldsymbol{u}}_{0})\in Y_{1}$ and (unique) ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$.
Let $r_{1}\geq\max\{r_{0},r^{\prime}_{1}\}$, (Lemmas 5.17 and 5.18). The proof now proceeds by showing that we can find $z\otimes z^{\prime}$ and ${\boldsymbol{t}}_{0}$ from $x[{\mathcal{V}}^{r_{1}}]$.
Since ${\mathcal{V}}$ is not perpendicular to ${\boldsymbol{v}}_{0}$ or ${\boldsymbol{v}}_{1}$, it crosses every 00-grid line $\ell_{0,n_{1}}$ and every $1$-grid line $\ell_{1,n_{0}}$ in $y({\boldsymbol{u}})$. Thus by Lemma 5.11, ${\mathcal{V}}$ crosses every bent 00-grid line $\lambda_{0,n_{1}}$ and every bent $1$-grid line $\lambda_{1,n_{0}}$ in $x$.
For each $n_{0}\in\mathbb{Z}$, we use (5.5) to find ${\boldsymbol{s}}_{0}\in{\mathcal{V}}\cap\lambda_{1,n_{0}}$ and $n^{\prime}_{1}\in\mathbb{Z}$ so that $w_{(n_{0},n^{\prime}_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}}_{0})\subseteq{\mathcal{V}}^{r_{1}}$. Similarly, we use (5.6) to find ${\boldsymbol{s}}_{1}\in{\mathcal{V}}\cap\lambda_{0,n_{1}}(x)$ and $n^{\prime}_{0}\in\mathbb{Z}$ so that $w_{(n^{\prime}_{0},n_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}}_{1})\subseteq{\mathcal{V}}^{r_{1}}$. Let $J:=\{{\boldsymbol{n}}\in\mathbb{Z}^{2}:w_{{\boldsymbol{n}}}(x)\subseteq{\mathcal{V}}^{r_{1}}\}$. Since ${\mathcal{V}}^{r_{1}}=\bigcup_{{\boldsymbol{s}}\in{\mathcal{V}}}B_{r_{1}}({\boldsymbol{s}})$, we see that $\pi_{0}(J)=\pi_{1}(J)=\mathbb{Z}$. Thus $x[{\mathcal{V}}^{r_{1}}]$ determines $(z\otimes z^{\prime})[J]$, which by Lemma 2.6 determines $z\otimes z^{\prime}$.
Now suppose $x^{\prime}=y({\boldsymbol{u}}_{0})^{*}$ from above, so that $x=T^{{\boldsymbol{t}}_{0}}x^{\prime}$ with $||{\boldsymbol{t}}_{0}||<4/3$. Let ${\boldsymbol{t}}^{\prime}_{0}\in\mathbb{R}^{2}$ satisfy $||{\boldsymbol{t}}^{\prime}_{0}||<4/3$ and suppose $x[{\mathcal{V}}^{r_{1}}]=(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[{\mathcal{V}}^{r_{1}}]$, (noting this holds for ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$). Then since $r_{1}\geq r_{0}$, we have $(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]$ and Lemma 5.17 implies ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$. ∎
Proof of Proposition 5.16.
The diameter of the regular decagon $D$ with edge length $2/5$ is $r^{\prime}_{2}=(2/5)(1+\sqrt{5})$. If $D$ sits in the upper half plane with an edge on the $x$-axis, centered along the $y$-axis, then $D\cup(-D)\subseteq B_{r^{\prime}_{2}}({{\boldsymbol{0}}})$. The same is true if one or both of the decagons is replaced with one or both of the hexagons $H_{1}$ and $H_{2}$ from the unfilled worms (in the upright position shown in Figure 6, (b)). It follows that if ${\boldsymbol{s}}$ lies in the central axis of a worm or cartwheel in a singular Penrose tiling $x$, then $B_{r^{\prime}_{2}}({\boldsymbol{s}})$ contains either an entire tiled hexagon or tiled decagon.
Now let $x$ be a singular Penrose tiling. Assume $r_{2}\geq\max\{r_{1},r^{\prime}_{2}\}$. We can apply the proof of Proposition 5.15, noting that ${\mathcal{V}}$ is not perpendicular to ${\boldsymbol{v}}_{0}$ or ${\boldsymbol{v}}_{1}$, to see that $x[{\mathcal{V}}^{r_{2}}]$ determines a tiling $y({\boldsymbol{u}})^{*}$ which becomes $x$ after filling either (a) an unfilled worm or (b) an unfilled cartwheel. Thus it remains to show that the direction of filling the worm ($+$ or --) or the cartwheel ($0,1,\dots,8$ or $9$) can be determined from $x[{\mathcal{V}}^{r_{2}}]$.
By assumption, ${\mathcal{V}}$ is not perpendicular to any 5th root of unity ${\boldsymbol{v}}_{j}$, whereas the central axis $l$ of a worm, or a cartwheel spoke $l_{j}$ is perpendicular to some ${\boldsymbol{v}}_{j}$. In case (a) there is a single worm along an axis $l$. Let ${\boldsymbol{s}}$ be the point where $l$ and ${\mathcal{V}}$ intersect. Then by the choice of $r_{2}$ there is at least one hexagon $H$ in the unfilled worm in $y({\boldsymbol{u}})^{*}$ satisfying $H\subseteq B_{r_{1}}({\boldsymbol{s}})\subseteq{\mathcal{V}}^{r_{2}}$. That worm is filled in $x[{\mathcal{V}}^{r_{2}}]$, and the direction of that filling determines how the worm is filled in $x$.
Case (b) is the case that $x$ has a cartwheel. There are two sub-cases: (i) $B_{r_{2}}({\boldsymbol{s}})\subseteq{\mathcal{V}}^{r_{2}}$ contains an unfilled decagon in $y({\boldsymbol{u}})^{*}$. This decagon is filled in $x$, and that filling determines the direction of filling for all of the spokes. (ii) ${\mathcal{V}}$ crosses all five spokes, $l_{j}$, $j=0,1,\dots,4$. If ${\boldsymbol{s}}_{j}$ is the intersection of $l_{j}$ with ${\mathcal{V}}$, then there is a hexagon $H^{j}\subseteq B_{r_{2}}({\boldsymbol{s}}_{j})\subseteq{\mathcal{V}}^{r_{2}}$ in $y({\boldsymbol{u}})^{*}$. The filling of each $H^{j}$ determines the filling of the spoke it is in, which, in turn, determines the filling of the decagon. ∎
6. The Penrose Wang tiles
6.1. Wang tiles
For $n,m>1$, let ${\mathcal{A}}=\{0,\dots,n-1\}$ be a finite alphabet, and let ${\mathcal{C}}_{{\rm b},{\rm t}}=\{0,..,m_{{\rm b},{\rm t}}-1\}$ and ${\mathcal{C}}_{{\rm l},{\rm r}}=\{0,..,m_{{\rm l},{\rm r}}-1\}$ be two finite sets of colors. Formally, a Wang tile set is $q:=\{Q^{a}:a\in{\mathcal{A}}\}$ where, formally, each Wang tile $Q^{a}$ is a tuple $Q^{a}=(a,c^{a}_{\rm b},c^{a}_{\rm t},c^{a}_{\rm l},c^{a}_{\rm r})$ with $c^{a}_{\rm b},c^{a}_{\rm t}\in{\mathcal{C}}_{{\rm b},{\rm t}}$ and $c^{a}_{\rm l},c^{a}_{\rm r}\in{\mathcal{C}}_{{\rm l},{\rm r}}$. More concretely, we view the Wang tile $Q^{a}$ as a unit square tile (shape ${\boldsymbol{e}}_{0}\wedge{\boldsymbol{e}}_{1}$) with the label444 Putting labels on a Wang tiles is not conventional, but it allows two different Wang tiles to have the same edge colors. This makes it possible to realize the $\mathbb{Z}^{2}$ full shift with Wang tiles. $a$, and colored edges $c^{a}_{\rm b},c^{a}_{\rm t},c^{a}_{\rm l},c^{a}_{\rm r}$: bottom, top, left and right. We will interpret a Wang tile set $q$ as an FLC prototile set by defining $q^{(2)}$ to be the condition that the tiles must meet edge to edge, and the colors of adjacent edges must match, and we let $W_{q}$ be the corresponding FLC tiling space. If $W_{q}\neq\emptyset$ (see Section 6.2), then $(W_{q},\mathbb{R}^{2},T)$ is an FLC tiling dynamical system that we call a Wang tile dynamical system.
Alternatively, we can interpret a Wang tile set $q$ as a recipe for $1$-step $\mathbb{Z}^{2}$ SFT $(Z_{q},\mathbb{Z}^{2},S)$, called a Wang tile $\mathbb{Z}^{2}$ SFT. For this, we define $Z_{q}\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$ in terms of horizontal (and vertical) permitted words ${\mathcal{P}}_{0}$ (and ${\mathcal{P}}_{1}$) where $a\,a^{\prime}\in{\mathcal{P}}_{0}$ if and only if $c^{a}_{\rm r}=c^{a^{\prime}}_{\rm l}$ (and $(a\,a^{\prime})^{t}\in{\mathcal{P}}_{1}$ (i.e., $a^{\prime}$ below $a$), if and only if $c^{a}_{\rm b}=c^{a^{\prime}}_{\rm t}$).
Remark 6.1.
There is a 1:1 correspondence between $Z_{q}\times[0,1)^{2}$ and $W_{q}$, in which $(z,{\boldsymbol{u}}^{\prime})$ with $z=(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q}$, and ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$ corresponds to $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}+{\boldsymbol{u}}^{\prime}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in W_{q}$.
Now suppose that $(Z,\mathbb{Z}^{2},S)$ is a $1$-step $\mathbb{Z}^{2}$ SFT, $Z\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$, defined by the horizontal and vertical permitted words ${\mathcal{P}}_{0}$ and ${\mathcal{P}}_{1}$ (see Section 2.3). We say $(Z,\mathbb{Z}^{2},S)$ is colorable if for some finite sets ${\mathcal{C}}_{{\rm b},{\rm t}}$ and ${\mathcal{C}}_{{\rm l},{\rm r}}$, there are functions (i) $\xi_{\rm l},\xi_{\rm r}:{\mathcal{A}}\to{\mathcal{C}}_{{\rm l},{\rm r}}$ so that $aa^{\prime}\in{\mathcal{P}}_{0}$ if and only if $\xi_{\rm r}(a)=\xi_{\rm l}(a^{\prime})$, and (ii) $\xi_{\rm b},\xi_{\rm t}:{\mathcal{A}}\to{\mathcal{C}}_{{\rm b},{\rm t}}$ so that $(aa^{\prime})^{t}\in{\mathcal{P}}_{1}$ if and only if $\xi_{\rm b}(a)=\xi_{\rm t}(a^{\prime})$. Clearly, if $(Z_{q},\mathbb{Z}^{2},S)$ is the $1$-step $\mathbb{Z}^{2}$ SFT that comes from a Wang tile set $q$, then it is colorable. Conversely, suppose that $(Z,\mathbb{Z}^{2},S)$ is a colorable $1$-step SFT. For each $a\in{\mathcal{A}}$ define the Wang tile $Q^{a}:=(a,\xi_{\rm b}(a),\xi_{\rm t}(a),\xi_{\rm l}(a),\xi_{\rm r}(a))\in q$. Then $(Z_{q},\mathbb{Z}^{2},S)$ is identified with $(Z,\mathbb{Z}^{2},S)$ by the map that reads symbols. Note that our formulation of Wang tiles here allows for $Z_{q}={\mathcal{A}}^{\mathbb{Z}^{2}}$ to be a full shift.
Proposition 6.2.
Let $(W_{q},\mathbb{R}^{2},T)$ be a Wang tile dynamical system and let $(Z_{q},\mathbb{Z}^{2},S)$ be the corresponding $\mathbb{Z}^{2}$ SFT. Then a direction ${\mathcal{V}}\subseteq\mathbb{R}^{2}$ is expansive for $(W_{q},\mathbb{Z}^{2},T)$ if and only if ${\mathcal{V}}$ is expansive for $(Z_{q},\mathbb{Z}^{2},S)$. The same result holds for any $\mathbb{Z}^{2}$ subshift $Z\subseteq Z_{q}$, and the corresponding tiling space $W\subseteq W_{q}$.
Proof.
This follows from Lemma 2.5, Definition 3.12, Proposition 3.13, and Theorem 3.10. ∎
6.2. Brief history
Wang tiles were first described by Hao Wang in 1961 [37], who posed the question, now known as Wang’s problem: is it possible to determine algorithmically (from $q$ and $q^{(2)}$) whether or not $W_{q}$ is nonempty? Wang proved (see [11]) that the problem is solvable in the class of Wang tile sets $q$ such that any $W_{q}\neq\emptyset$ contains a doubly periodic tiling $x$ (i.e., $T^{n_{0}{\boldsymbol{e}}_{0}}x=x$ and $T^{n_{1}{\boldsymbol{e}}_{1}}x=x$ for some $n_{0},n_{1}>0$), and he conjectured that all Wang tile sets $q$ belong to this class. Soon thereafter, however, Berger [2] showed that there exists a Wang tile set $q$ with $W_{q}\neq\emptyset$ such that every $x\in W_{q}$ is aperiodic. In such a case, it is customary to say $q$ is an aperiodic Wang tile set. Moreover, Berger [2] showed that in general, Wang’s problem is undecidable.
In terms of dynamics, aperiodicity for the Wang tile set $q$ means that $(W_{q},\mathbb{R}^{2},T)$ is free. Berger’s [2] apreiodic set $q$ of Wang tiles has #$\#(q)>20,\!000$, which led to a race to reduce this number (see [11] for a nice account). Recently, Jeandel and Rao [18] ended this race by constructing an aperiodic Wang tile set $q_{\rm jr}$ with #$\#(q_{\rm jr})=11$ and also showing that #$\#(q_{\rm jr})=11$ is the smallest possible size.
Lebbé showed [20] that the Jaendel/Rao SFT $(Z_{\!q_{\rm jr}},\mathbb{Z}^{2},S)$ is uniquely ergodic but not minimal [20]. However, $(Z_{\!q_{\rm jr}},\mathbb{Z}^{2},S)$ has a unique minimal subshift $(Z_{0},\mathbb{Z}^{2},S)$, and that subshift has $\mathbb{N}_{1}(S)=\{0,\gamma+3,-3\gamma+2,-\gamma+5/2\}$, [21], (but a priori, the set $Z_{\!q_{\rm jr}}\backslash Z_{0}$ could be responsible for additional non-expansive directions). Finally, Lebbé showed that $(Z_{0},\mathbb{Z}^{2},S)$ is almost automorphic, with $\Sigma_{S}=(\mathbb{Z}[\alpha]/\mathbb{Z})\times(\mathbb{Z}[\alpha]/\mathbb{Z})$, the same as Raphael Robinson’s Penrose Wang tile SFT, (Section 6.4). However, the two are not topologically conjugate.
6.3. Suspensions, compositions, deformations, and continuous orbit equivalence
Given a dynamical system $(Z,\mathbb{Z}^{2},S)$, let ${\widetilde{Z}}=Z\times[0,1]^{2}/\sim$, where $(z,(1,s_{1}))\sim(S^{{\boldsymbol{e}}_{0}}z,(0,s_{1}))$ and $(z,(s_{0},1))\sim(S^{{\boldsymbol{e}}_{1}}z,(s_{0},0))$ for $z\in Z$ and ${\boldsymbol{s}}:=(s_{0},s_{1})\in[0,1]^{2}$. The unit suspension of $(Z,\mathbb{Z}^{2},S)$ is the dynamical system $({\widetilde{Z}},\mathbb{R}^{2},{\widetilde{S}})$ defined by ${\widetilde{S}}^{\boldsymbol{t}}(z,{\boldsymbol{s}})=(S^{\lfloor{\boldsymbol{t}}\rfloor}z,\{{\boldsymbol{t}}+{\boldsymbol{s}}\})$. For a Wang tiling system $(W_{q},\mathbb{R}^{2},T)$, where $q=\{Q^{j}:j\in{\mathcal{A}}\}$, and $(Z_{q},\mathbb{Z}^{2},S)$ is the corresponding SFT, clearly $(\widetilde{Z_{q}},\mathbb{R}^{2},\widetilde{S})$ is topologically conjugate to $(W_{q},\mathbb{R}^{2},T)$.
Next, we change the shape of the Wang tiles $q$ from squares to a parallelograms. Let ${\boldsymbol{f}}_{0},{\boldsymbol{f}}_{1}\in\mathbb{R}^{2}$ with $0<\angle{\boldsymbol{f}}_{0}{\boldsymbol{f}}_{1}<\pi$, and let $A=[{\boldsymbol{f}}_{0};{\boldsymbol{f}}_{1}]^{t}$. The parallelogram Wang tiles $Aq:=\{AQ^{a}:Q^{a}\in q\}$ are defined to be the copies of ${\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1}=A({\boldsymbol{e}}_{0}\wedge{\boldsymbol{e}}_{1})$ marked the same way as the squares in $q$. Following Remark 6.1, we write $v\in W_{Aq}$ by $v=\{AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$, where ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$ and $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q}$. Here $(Aq)^{(2)}:=A(q^{(2)})$.
A parallelogram Wang tiling dynamical system $(W_{Aq},\mathbb{R}^{2},T)$ is no longer the unit suspension of $(Z_{q},\mathbb{Z}^{2},S)$. However, we can recover it from $({\widetilde{Z_{q}}},\mathbb{R}^{2},{\widetilde{S}})$ as follows: For an $\mathbb{R}^{2}$-action $(X,\mathbb{R}^{2},T)$, and a matrix $A\in Gl(2,\mathbb{R})$, we define the composition action by $(T\circ A)^{\boldsymbol{t}}x:=T^{A{\boldsymbol{t}}}x$.
Lemma 6.3.
  1. (1)
    The parallelogram Wang tile system $(W_{Aq},\mathbb{R}^{2},T)$ is topologically conjugate to the composition action $({\widetilde{Z_{q}}},\mathbb{R}^{2},{\widetilde{S}}\circ A)$.
  2. (2)
    The parallelogram Wang tile system $(W_{Aq},\mathbb{R}^{2},T)$ is continuously orbit equivalent (Definition 2.1) to the Wang tile system $(W_{q},\mathbb{R}^{2},T)$.
  3. (3)
    The (weakly and strongly) expansive directions for $(W_{Aq},\mathbb{R}^{2},T)$ are the directions $A({\mathcal{V}})$ for the expansive directions ${\mathcal{V}}$ of $(Z_{q},\mathbb{Z}^{2},S)$.
Proof.
For (1), we construct a model for ${\widetilde{S}}\circ A$ on $Z_{q}\times({\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1})$. The conjugacy is the homeomorphism ${\rm id}\times A:Z_{q}\times[0,1)^{2}\to Z_{q}\times({\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1})$. Part (2) follows from part (1), and the fact $(W_{q},\mathbb{R}^{2},T)$ and $(W_{q},\mathbb{R}^{2},T\circ A)$, (which is topologically conjugate to $(W_{Aq},\mathbb{R}^{2},T)$), both act on $W_{q}$, and have the same orbits. Part (3) follows from Propositions 3.13 and 3.10. ∎
Remarks 6.4.
(See e.g. [32]):
  1. (1)
    Suppose that $\widetilde{T}$ is the unit suspension of $T$. Then $\Sigma_{\widetilde{T}}=\pi^{-1}\Sigma_{T}$ where $\pi({\boldsymbol{s}}):=\{{\boldsymbol{s}}\}:\mathbb{R}^{2}\to\mathbb{T}^{2}$.
  2. (2)
    The point spectrum of an $\mathbb{R}^{2}$ composition action $T\circ A$ satisfies $\Sigma_{T\circ A}=A^{t}\Sigma_{T}$.
Remark 6.5.
In the terminology of [35] and [5], a deformation $\psi$ is a continuous orbit equivalence (i.e., orbit preserving homeomorphism) between two FLC tiling spaces such that the restriction to each tile depends only on its prototile type. The point of view in [35] is closer to ours here [^5] The point of view in [5] is in terms of the Anderson-Putnam complex., and in particular, it is shown in [35] that any $\mathbb{R}^{2}$ FLC tiling space $(X,\mathbb{R}^{2},T)$ is continuously orbit equivalent to the unit suspension of $\mathbb{Z}^{2}$ subshift, essentially realized as a subspace of a square Wang tiling space (any full shift can be realized by Wang tiles). Some examples of deformation are discussed in Section 6.4.
6.4. The Penrose Wang tiles
From now on, ${\mathcal{A}}:=\{0,\dots,23\}$. In Section 5.3 we discussed the $24$ Wang patches $\{w^{a}:a\in{\mathcal{A}}\}$, shown in Figure 9. Here we will construct the corresponding $24$ Penrose Wang tiles due to Raphael Robinson ([11], Exercise 11.1.2), made by deforming the Wang patches, as well as some related tiles and tilings.
For each the bottom and top margin ($s\in\{{\rm b},{\rm t}\}$) of each Wang patch $w^{a}$, we assign a finite word edge code ${\beta}^{a}_{s}\in\{\overline{2},\overline{3},4\}^{*}$ that records the tile edges along the outside of that margin. Similarly, for each left and right margin ($s\in\{{\rm l},{\rm r}\}$) we assign an edge code ${\beta}^{a}_{s}\in\{2,\overline{3},\overline{4}\}^{*}$. We record, $j$ for ${\boldsymbol{v}}^{\prime}_{j}$ and $\overline{j}$ for $-{\boldsymbol{v}}^{\prime}_{j}$, skipping the corner tiles ($j=0$ or $j=1$) at either end. The edges are read left to right for $\beta_{\rm b}^{a}$ and $\beta_{\rm t}^{a}$, and up for $\beta_{\rm l}^{a}$ and $\beta_{\rm r}^{a}$. For example, since the edges along the bottom margin in $w^{0}$ are ${\boldsymbol{v}}^{\prime}_{0},{\boldsymbol{v}}^{\prime}_{4},-{\boldsymbol{v}}^{\prime}_{3},-{\boldsymbol{v}}^{\prime}_{2},{\boldsymbol{v}}^{\prime}_{0}$, we record $\beta^{0}_{\rm b}=4\overline{3}\overline{2}$. The edge codes are shown, for each $a\in{\mathcal{A}}$, in the second column of Table 1.
$a\in{\mathcal{A}}$ $\beta^{a}_{\rm b},\beta^{a}_{\rm t},\text{ and }\beta^{a}_{\rm l},\beta^{a}_{\rm r}$ $c^{a}_{\rm b},c^{a}_{\rm t}\text{ and }c^{a}_{\rm l},c^{a}_{\rm r}$ $v\text{ for }{\boldsymbol{b}}^{v}_{\rm b},{\boldsymbol{b}}^{v}_{\rm t}\text{ and }{\boldsymbol{b}}^{v}_{\rm l},{\boldsymbol{b}}^{v}_{\rm r}$ $t\text{ for }U^{t}$
0 43¯2¯,42¯, $4\overline{3}\overline{2},4\overline{2},\ 2\overline{3}\overline{4},2\overline{4}$, 0,1, 0,1 0,1, 0,1 0
1 $2{2\overline{4},2\overline{3}\overline{4}},\ \overline{4}2,\overline{4}\overline{3}$ 2 2,3, 2,3 1,0, 1,0 1
2 $\overline{2},\overline{3}\overline{2}\ ,\overline{4},\overline{3}\overline{4}$ 4,5, 4,5 2,3, 2,3 2
3 $\overline{2}\overline{3},\overline{2},\ \overline{4}\overline{3},\overline{4}$ 6,4, 6,4 3,2, 3,2 3
4 $\overline{3}\overline{2}4,\overline{2}\overline{3}4,\ \overline{3}\overline{4}2,\overline{4}\overline{3}2$ 7,3, 7,3 0,0, 0,0 4
5 43¯2¯,42¯3¯, $4\overline{3}\overline{2},4\overline{2}\overline{3},\ 2\overline{3}\overline{4},2\overline{4}\overline{3}$ 0,8, 0,8 0,0, 0,0 4
6 42¯,3¯2¯4, $24\overline{2},\overline{3}\overline{2}4,\ 2\overline{4},\overline{3}\overline{4}$ 2 1,7, 1,7 1,0, 1,0 1
7 42¯3¯,2¯4, $24\overline{2}\overline{3},\overline{2}4,\ 2\overline{4}\overline{3},\overline{4}$ 2 8,2, 8,2 0,1, 0,1 0
8 $\overline{2}\overline{3}4,4\overline{2}\overline{3},\ \overline{3}\overline{4},\overline{4}\overline{3}$ 3,8, 5,6 0,0, 3,3 5
9 ${\overline{3}\overline{2}4},4\overline{3}\overline{2},\ \overline{3}\overline{4},\overline{4}\overline{3}$ 7,0, 5,6 0,0, 3,3 5
10 $\overline{3}\overline{2},\overline{2}\overline{3},\ \overline{4}\overline{3}2,2\overline{4}\overline{3}$ 5,6, 3,8 3,3, 0,0 6
11 $\overline{3}\overline{2},\overline{2}\overline{3},\ \overline{3}\overline{4}2,2\overline{3}\overline{4}$ 5,6, 7,0 3,3, 0,0 6
12 $\overline{2}\overline{3},\overline{2},\ \overline{4}\overline{3}2,\overline{4}2$ 6,4, 3,2 3,2, 0,1 7
13 2¯,3¯2¯, $\overline{2},\overline{3}\overline{2},\ 2\overline{4},\overline{4}2\overline{3}$ 4,5, 1,0 2,3, 1,0 8
14 $\overline{2}\overline{3}4,\overline{2}4,\ ,\overline{4}\overline{3},\overline{4}$ 3,2, 6,4 0,1, 3,2 9
15 $4\overline{2},4\overline{3}\overline{2},\ \overline{4},\overline{3}\overline{4}$ 1,0, 4,5 1,0, 2,3 10
16 2¯4,3¯2¯4, $\overline{2}4,\overline{3}\overline{2}4,\ 2\overline{4},\overline{4}\overline{3}2$ 2,7, 1,3 1,0, 1,0 1
17 42¯3¯,42¯, $24\overline{2}\overline{3},4\overline{2},\ 2\overline{3}\overline{4},\overline{4}$ 2 8,1, 0,2 0,1, 0,1 0
18 43¯2¯,2¯4, $44\overline{3}\overline{2},\overline{2}4,\ 2\overline{4}\overline{3},\overline{2}$ 4 0,2, 8,1 0,1, 0,1 0
19 $24\overline{2},\overline{2}\overline{3}4,\ \overline{4}2,\overline{3}\overline{4}$ 2 1,3, 2,7 1,0, 1,0 1
20 $\overline{2},\overline{2}\overline{3},\ \overline{4}2,2\overline{3}\overline{4}$ 4,6, 2,0 2,3, 1,0 8
21 $\overline{3}\overline{2},\overline{2},\ \overline{4}\overline{3}2,2\overline{4}$ 5,4, 3,1 3,2, 0,1 7
22 $\overline{2}4,4\overline{3}\overline{2},\ \overline{4},\overline{4}\overline{3}$ 2,0, 4,6 1,0, 2,3 10
23 $\overline{2}\overline{3}4,4\overline{2},\ \overline{3}\overline{4},\overline{4}$ 3,1, 5,4 0,1, 3,2 9
Table 1. Columns 1-3: edge codes for Wang patches and corresponding Wang tile edge colors. Columns 4 and 5: edge vector numbers for tetragon Wang tiles, and tetragon Wang tile type.
Column 2 of Table 1 shows that there are nine distinct top and bottom code words, and for each of them we assign a color $c_{\rm b},c_{\rm t}\in{\mathcal{C}}_{{\rm b},{\rm t}}=\{0,1,\dots,8\}$. Similarly, there are nine distinct left and right code words each assigned a color $c_{\rm l},c_{\rm r}\in{\mathcal{C}}_{{\rm l},{\rm r}}=\{0,1,\dots,8\}$. The colors are shown in Table 1, Column 3.
Definition 6.6.
For each row in Table 1, (each $a\in{\mathcal{A}}$), we define a Penrose Wang tile $Q^{a}:=(a,c_{\rm b},c_{\rm t},c_{\rm l},c_{\rm r})$, where $c_{\rm b},c_{\rm t},c_{\rm l},c_{\rm r}$ are the colors shown in the Column 3. The $24$ Penrose Wang tiles are given by $q_{\rm p}=\{Q^{a}:a\in{\mathcal{A}}\}$, where $q_{\rm p}^{(2)}$ is as described for Wang tiles in general in Section 6.1.
Two tiles $Q^{a}$ and $Q^{a^{\prime}}$ can be vertically (or horizontally) adjacent in some Wang tiling in $W_{q_{\rm p}}$ if and only if the same is true for the corresponding Wang patches $w^{a}$ and $w^{a^{\prime}}$ for some $x\in\mathsf{X}$. The Penrose Wang tile dynamical system is denoted by $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, and the corresponding $\mathbb{Z}^{2}$ SFT is denoted by $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$, where we write $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q_{\rm p}}$.
Corollary 6.7.
Let $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}+{\boldsymbol{u}}^{\prime}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in W_{q_{\rm p}}$, ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$, $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q_{\rm p}}$, be an arbitrary Penrose Wang tiling, ((((see Remark 6.1)))). Then there is a Penrose tiling $x\in\mathsf{X}$, unique up to translation, so that $w_{\boldsymbol{n}}(x)$ is type $w^{a_{\boldsymbol{n}}}$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Conversely, for each Penrose tiling $x\in\mathsf{X}$ there is a Penrose Wang tiling $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$, unique up to translation, so that $w_{\boldsymbol{n}}(x)$ is type $w^{a_{\boldsymbol{n}}}$ for each ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. In particular, $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(\mathsf{X},\mathbb{R}^{2},T)$ are continuously orbit equivalent (via a deformation), and $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ are nonempty and free.
Definition 6.8.
We define Penrose parallelogram Wang tiles as follows: For ${\boldsymbol{f}}_{0}$ and ${\boldsymbol{f}}_{1}$ in (5.2), let $A=[{\boldsymbol{f}}_{1};{\boldsymbol{f}}_{0}]^{t}$, so that $A[0,1]^{2}={\boldsymbol{f}}_{1}\wedge{\boldsymbol{f}}_{0}=R$, the rhombus that supports the grid patches. The Penrose parallelogram Wang tiles are the tiles $Aq_{\rm p}=\{AQ^{a}:a\in{\mathcal{A}}\}$, which differ from the square Penrose Wang tiles $q_{\rm p}$ only in their shape.
Like $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, the Penrose parallelogram Wang tile dynamical system $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$ is nonempty, free, and is Kakutanni equivalent to $(\mathsf{X},\mathbb{R}^{2},T)$. More of its properties are described in Theorem 6.11 below.
Next, we recall the unmarked Penrose tetragon tilings $u=\{U_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$, described in Definition 5.12. For any $x\in\mathsf{X}$, the vertices are ${\boldsymbol{d}}_{\boldsymbol{n}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$ and ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$. We denote the vectors along the four sides by
${\boldsymbol{b}}_{\rm b}$ $:={\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}-{\boldsymbol{d}}_{\boldsymbol{n}},$ ${\boldsymbol{b}}_{\rm t}:={\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}-{\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}},$
${\boldsymbol{b}}_{\rm l}$ $:={\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}-{\boldsymbol{d}}_{\boldsymbol{n}},\text{\ \ \ \ \ \ \ \ \ \ and }$ ${\boldsymbol{b}}_{\rm r}:={\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}-{\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}.$
These vectors can also be found by summing the edge vectors along the sides of the corresponding Wang patch $w_{\boldsymbol{n}}(x)$:
Lemma 6.9.
Suppose that $U_{\boldsymbol{n}}(x)$ is the unmarked tetragon Penrose tile corresponding to the Wang patch $w_{\boldsymbol{n}}(x)$ in some $x\in\mathsf{X}$ ((((we think of $U_{\boldsymbol{n}}$ as type $U^{a}$ if the type of $w_{\boldsymbol{n}}(x)$ is $w^{a}$)))). Then for each $s\in\{{\rm b},{\rm t}\}$ and $s^{\prime}\in\{{\rm l},{\rm r}\}$,
${\boldsymbol{b}}_{s}^{a}={\boldsymbol{v}}^{\prime}_{0}+\sum_{j\in\beta_{s}^{a}}{\boldsymbol{v}}^{\prime}_{j}\text{ and }{\boldsymbol{b}}_{s^{\prime}}^{a}={\boldsymbol{v}}^{\prime}_{1}+\sum_{j\in\beta_{s}^{a}}{\boldsymbol{v}}^{\prime}_{j}.$ (6.1)
In particular, ${\boldsymbol{b}}_{s}^{a}$ for $s\in\{{\rm b},{\rm t},{\rm l},{\rm r}\}$ depends only on the code word $\beta_{s}^{a}$, or equivalently, the color $c_{s}^{a}$ (Table 1, Column 2).
Computing the sums (6.1) for each $a\in{\mathcal{A}}$ shows that there are four possible vectors for ${\boldsymbol{b}}_{\rm b}^{a}$ and ${\boldsymbol{b}}_{\rm t}^{a}$, and four more for ${\boldsymbol{b}}_{\rm l}^{a}$ and ${\boldsymbol{b}}_{\rm r}^{a}$. These are shown in Columns 2 and 5 on Table 2 as ${\boldsymbol{b}}_{\rm b}^{v},{\boldsymbol{b}}_{\rm t}^{v}$ for $v\in\{0,1,2,3\}$, and as ${\boldsymbol{b}}_{\rm l}^{v^{\prime}},{\boldsymbol{b}}_{\rm r}^{v^{\prime}}$ for $v^{\prime}\in\{0,1,2,3\}$. Columns 3 and 6 of Table 2 show all the (Wang tile) colors $c_{\rm b},c_{\rm t},c_{\rm l}$, or $c_{\rm r}$ that correspond to these vectors, as described in Lemma 6.9.
Figure 13. The 24 tetragon Penrose tiles $p_{4}:=\{M^{a}:a\in{\mathcal{A}}\}$, where $a$ is shown in the rounded squares. Each edge is labeled with its color, $c_{\rm b},c_{\rm t}\in\{{\it 0},\dots\it 8\}$ and $c_{\rm l},c_{\rm r}\in\{{\it 0},\dots\it 8\}$, (Table 1, Column 3), with the subscript showing the side vector $v$ or $v^{\prime}$ (Table 2, Columns 2 and 4). The tile shape (the unmarked tetragon tile type $v$) is shown in the circle below each tile. The 24 Penrose Wang tiles (or Penrose parallelogram Wang tiles) may be obtained by deforming each tetragon with the unit square or rhomb.
Column 4 of Table 1 shows which edge vector defines each of the four sides of the unmarked tetragon Penrose tile. Table 1, Column 5 shows that there are 11 different unmarked tetragon Penrose tiles $U^{t}$ for $t\in\{0,\dots,10\}$.
$v$ ${\boldsymbol{b}}^{v}_{\rm b},{\boldsymbol{b}}^{v}_{\rm t}$ $c_{\rm b},c_{\rm t}$ $v^{\prime}$ ${\boldsymbol{b}}^{v^{\prime}}_{\rm l},{\boldsymbol{b}}^{v^{\prime}}_{\rm r}$ $c_{\rm l},c_{\rm r}$
0 $(1.1710,-0.3804)$ 0,3,7,8 0 $(0,1.231)$ 0,3,7,8
1 $(0.8472,-0.6155)$ 1,2 1 $(-0.3236,0.9960)$ 1,2
2 $(0.7236,-0.2351)$ 4 2 $(0,0.7608)$ 4
3 $(1.047,0)$ 5,6 3 $(0.3236,0.9960)$ 5,6
Table 2. The vertical and horizontal edge vectors for marked and unmarked tetragon Penrose tiles. The vectors for $v=0$ and $v=2$, and for $v^{\prime}=0$ and $v^{\prime}=2$, have the same slopes. Columns 3 and 5 show which edge colors correspond to each vector.
Definition 6.10.
We define marked tetragon Penrose tiles $m_{4}=\{M^{a},a\in{\mathcal{A}}\}$ as follows. For each $a\in{\mathcal{A}}$ (Table 1, Column 1) we define $M^{a}$ by decorating the corresponding unmarked tetragon Penrose tile $U^{v}$, where $v=v(a)$ is given in Table 1, Column 5, by coloring its edges using the Wang tile colors $c_{\rm b},c_{\rm t},c_{\rm l}$, and $c_{\rm r}$ (Table 1, Column 3). The matching rule $m_{4}^{(2)}$ is the requirement that the colors of the adjacent edges of two adjacent tiles must match.
Any two horizontal or vertical edges that have the same color also have the same edge vector. Thus the adjacency condition is the same as for the Penrose Wang tiles. We call $(X_{m_{4}},\mathbb{R}^{2},T)$ the tetragon Penrose tiling dynamical system. We see that $(X_{m_{4}},\mathbb{R}^{2},T)$ is continuously orbit equivalent to $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ by a deformation, so in particular, $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ is nonempty and free.
The next two results summarize the dynamical properties of the four dynamical systems studied in this section. Theorem 6.11 discusses the tetragon Penrose tiling system $(X_{m_{4}},\mathbb{R}^{2},T)$ and Penrose parallelogram Wang tile system $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$. Corollary 6.12 discusses Raphael Robinson’s Penrose Wang tile system $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, and the corresponding $\mathbb{Z}^{2}$ SFT $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$.
Theorem 6.11.
The dynamical systems $(X_{m_{4}},\mathbb{R}^{2},T)$ and $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$ are both topologically conjugate ((((MLD in the first case)))) to the Penrose tiling dynamical system $(\mathsf{X},\mathbb{R}^{2},T)$. Each is strictly ergodic and almost automorphic, with $\Sigma_{T}=Z[e^{2\pi i/5}]\subseteq\mathbb{R}^{2}$, $\mathfrak{M}_{T}=\{1,2,10\}$, and $\mathbb{N}_{1}(T)=\{{\boldsymbol{v}}_{0}^{\perp},\dots,{\boldsymbol{v}}_{4}^{\perp}\}$.
Corollary 6.12.
The dynamical systems $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ are both strictly ergodic, and almost automorphic, satisfying $\Sigma_{T}=\mathbb{Z}[\alpha]\times\mathbb{Z}[\alpha]\subseteq\mathbb{R}^{2}$ and $\Sigma_{S}=(\mathbb{Z}[\alpha]/\mathbb{Z})\times(\mathbb{Z}[\alpha]/\mathbb{Z})\subseteq\mathbb{T}^{2}$. Also $\mathfrak{M}_{T}=\mathfrak{M}_{S}=\{1,2,10\}$, and $\mathbb{N}_{1}(T)=\mathbb{N}_{1}(S)=\{0,1,\infty,\gamma,\alpha\}$.
Proof.
This follows from Theorem 6.11, Lemma 6.3, and Remark 6.4. However, $\Sigma_{S}=(\mathbb{Z}[\alpha]/\mathbb{Z})\times(\mathbb{Z}[\alpha]/\mathbb{Z})$ is easier to see from the fact that $R_{\alpha}\otimes R_{\alpha}$ is the Kronecker factor of $S$. ∎
Remark 6.13.
In [16], the minimality of $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ is stated as a conjecture, and an earlier preprint version of [16]. The preprint [16] also proves that $\mathbb{N}_{1}(T)$ is as stated above.
Proof of Theorem 6.11.
It suffices to prove the topological conjugacies, and then everything else follows from Theorems 4.3 and 1.1. For $(X_{m_{4}},\mathbb{R}^{2},T)$ there is an obvious MLD topological conjugacy.
[…]
Next, we consider a different formula for the restriction $\Phi|_{\mathsf{X}_{1}}$ of $\Phi$ to $\mathsf{X}_{1}$. For $x\in\mathsf{X}_{1}$, let $y({\boldsymbol{u}})=\varphi(x)\in Y_{1}$. By Lemma 5.8, $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Define $\theta:Y_{1}\to W_{Aq_{\rm p}}$ to be the map that replaces each grid patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})$ with the Wang tile $\Theta(g_{\boldsymbol{n}}({\boldsymbol{u}})):=AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})$, noting that both $g_{\boldsymbol{n}}({\boldsymbol{u}})$ and $AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})$ are supported on $R_{\boldsymbol{n}}$. In particular, $\theta$ is a local map in the sense of Definition 3.5, so it is uniformly continuous, and satisfies $\theta(T^{\boldsymbol{t}}y({\boldsymbol{u}}))=T^{\boldsymbol{t}}(\theta(y({\boldsymbol{u}})))$. Clearly $\theta(\varphi(x))=\Phi|_{\mathsf{X}_{1}}(x)$ (from the first part), so $\Phi$ is uniformly continuous on $\mathsf{X}_{1}$, and satisfies $\Phi|_{\mathsf{X}_{1}}(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi|_{\mathsf{X}_{1}}(x))$. It follows that $\Phi$ is the unique continuous extension of $\theta(\varphi(x))$, and thus satisfies $\Phi(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi(x))$.
It remains to show $\Phi$ is a bijection. Lemma 6.7 shows there is a 1:1 correspondence between $T$ orbits on $\mathsf{X}$ and $T$ orbits on $W_{Aq_{\rm p}}$. But since $T$ acts freely on both $\mathsf{X}$ and $W_{Aq_{\rm p}}$, the fact that $\Phi(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi(x))$ shows that $\Phi$ is bijective on each orbit as well. This implies $\Phi$ is a bijection. ∎